### v1

# Combine and conquer: model averaging for out-of-distribution forecasting

Abstract
Travel behaviour modellers have an increasingly diverse set of models at their disposal, ranging from traditional econometric structures to models from mathematical psychology and data-driven approaches from machine learning. A key question arises as to how well these different models perform in prediction, especially when considering trips of different characteristics from those used in estimation, i.e. out-of-distribution prediction, and whether better predictions can be obtained by combining insights from the different models. Across two case studies, we show that while data-driven approaches excel in predicting mode choice for trips within the distance bands used in estimation, beyond that range, the picture is fuzzy. To leverage the relative advantages of the different model families and capitalise on the notion that multiple ‘weak’ models can result in more robust models, we put forward the use of a model averaging approach that allocates weights to different model families as a function of the distance between the characteristics of the trip for which predictions are made, and those used in model estimation. Overall, we see that the model averaging approach gives larger weight to models with stronger behavioural or econometric underpinnings the more we move outside the interval of trip distances covered in estimation. Across both case studies, we show that our model averaging approach obtains improved performance both on the estimation and validation data, and crucially also when predicting mode choices for trips of distances outside the range used in estimation.
Keywords: choice modelling; forecasting; machine learning; mathematical psychology; mode choice; model averaging
1 Introduction
The travel behaviour modelling literature has focussed extensively on two sorts of models, namely models for inference (henceforth inference models) and models for forecasting (henceforth forecasting models). Inference models aim to understand current travel behaviour (e.g. to recover the value of travel time), while forecasting models aim to predict future travel behaviour in new settings (e.g. due to changes in transport policies, such as toll roads and fuel levies).
Inference models and forecasting models are evaluated differently by analysts. When building an inference model, an analyst is keen to ensure that the model generates behaviourally plausible insights into explanatory factors and their relative impacts. When building a forecasting model, the focus is much more on a model’s ability to generate accurate predictions of behaviour. Importantly, in forecasting, we predict the behaviour in new settings, meaning that the models have to generalise out-of-sample. As some of the settings we wish to make predictions for may be outside the range of level-of-service values that the models have been estimated on, it is important for forecasting models to also generalise out-of-distribution.
Travel behaviour analysts are increasingly interested in a more diverse set of models, covering econometric structures, models from mathematical psychology, and machine learning techniques. With the typical model evaluation approach focussing on the ability of models to replicate the behaviour in the empirical data as well as possible, it should then come as no surprise that researchers in the travel behaviour field are increasingly attracted by the comparatively good prediction performance of machine learning approaches (cf. Hagenauer and Helbich, 2017).
Of course, the entire purpose of building a forecasting model is not to replicate choices that have already been made, but to make predictions for new settings. Despite the growing interest in machine learning approaches, the dominant - although not necessarily evidenced-based - view is that forecasting models with a solid behavioural underpinning, such as Random Utility Maximisation (RUM) based discrete choice models, are better equipped to predict behaviour under new settings than are models with a weak or no behaviour underpinning, such as e.g. machine learning models (cf. van Cranenburgh et al., 2022).
However, what is currently less well understood is how to value and incorporate model performance, i.e. the model fit on the empirical data, in the process of developing forecasting models. What is clear is that good model performance is insufficient to establish that a model will make good out-of-distribution predictions. For that, the model must have captured underlying relationships that generalise beyond the observed data. The choice modelling literature has at times recognised that more advanced models that offer a better fit on the empirical data do not necessarily lead to better forecasts (see e.g. Fox et al., 2014). However, the question of how to develop forecasting models considering the model’s performance and behavioural underpinning is still open. On the one hand, the fact that model parameters are estimated by maximising the model performance on the empirical data, i.e. how well they replicate current choices, is somewhat at odds with the aim of forecasting, i.e. to generalise out-of-distribution. On the other hand, it is intuitive that the performance of a model on the empirical data still pertains to relevant information on its ability to generalise out-of-distribution.
The aim of this paper is twofold. First, we aim to shed light on the prediction performance of forecasting models with varying levels of behavioural underpinning as a function of the “distance” between the training data and forecasting scenarios for models. Second, we aim to develop an approach that improves prediction performance in forecasts through a model averaging approach that assigns different weights to different models depending on this “distance". In our work, we use trip distance as the key variable, but the approach works with any variable that goes out of distribution, not only physical distance.
We employ a neural network for the model averaging. This allows learning potential nonlinear relations between distance and model weights. Using the resulting model averaging structure, we aim not only to develop a flexible tool for combining models but also to gain rule-of-thumb insights for the conditions under which specific sorts of models perform best in terms of out-of-distribution forecasting.
The remainder of this paper is organised as follows. Section 2 presents our proposed model averaging approach. This is followed by the two case studies in Section 3, and conclusions in Section 4.
2 Methodology
As outlined in the introduction, the methodological novelty of the present paper is the use of model averaging for out-of-distribution prediction. In this section, we first discuss the general concept of model averaging before looking into the adaptation required for out-of-distribution prediction.
2.1 Generic model averaging approach
In model averaging (see e.g. Hancock et al., 2020), we first estimate $M$ different models on the data, with model $m$ giving a likelihood $L_{n,t,m}\left(\Omega_{m}\right)$ for the choice observed by decision maker $n$ in choice situation $t$, using a vector of estimated model parameters $\Omega_{m}$. Let us further define $L_{n,m}\left(\Omega_{m}\right)$ to be the likelihood for the entire sequence of $T_{n}$ choices observed for person $n$ according to model $m$.
Model averaging now combines the insights from these different models into one overall structure. The specification of the model averaging log-likelihood function depends on whether an analyst wishes the weights for individual models to be defined at the person level (i.e. $\pi_{n,m}$ for model $m$ for person $n$) or the observation level (i.e. $\pi_{n,t,m}$ for model $m$ for person $n$ in choice situation $t$). With the former, we would have:
$LL\left(\Theta\right)=\sum_{n=1}^{N}log\sum_{m=1}^{M}\pi_{n,m}L_{n,m}\left(\hat{\Omega}_{m}\right),$ (1)
while, with the latter, we would have:
$LL\left(\Theta\right)=\sum_{n=1}^{N}\sum_{t=1}^{T_{n}}log\sum_{m=1}^{M}\pi_{n,t,m}L_{n,t,m}\left(\hat{\Omega}_{m}\right).$ (2)
with $T_{n}$ choices per person. In this notation, $\Theta$ is a vector of parameters for the model averaging model, while $\hat{\Omega}_{m}$ are the maximum likelihood estimates (MLE) for the vector of parameters $\Omega_{m}$ for model $m$.
The weights are parameterised as being a function of the person ($z_{n}$) and observation ($z_{n,t}$) characteristics, with parameters $\gamma_{m}$ estimated to capture the influence of these characteristics. In particular, we would have:
$\pi_{n,m}\left(\Theta\right)=\frac{e^{\gamma_{m}^{\prime}z_{n}}}{\sum_{l=1}^{M}e^{\gamma_{l}^{\prime}z_{n}}},$ (3)
or
$\pi_{n,t,m}\left(\Theta\right)=\frac{e^{\gamma_{m}^{\prime}z_{n,t}}}{\sum_{l=1}^{M}e^{\gamma_{l}^{\prime}z_{n,t}}},$ (4)
where an appropriate normalisation is needed for $\gamma$, typically setting $\gamma_{m}=0$ for one model $m$.
A reader familiar with latent class models will realise that what is described above is, in essence, a sequential latent class structure. The sequential nature of the process involves first estimating the parameters for the $M$ individual models before then keeping those parameters fixed (at $\hat{\Omega}_{m}$ for model $m$) and estimating weights for each model, as in Equation 1 and 2. In maximising the model averaging likelihood, an analyst thus, in fact, does not require the model parameters for the individual models, only the probabilities of each observed choice.
In applying the model to predict choices out-of-sample, an analyst then again requires the predicted choice probabilities from the individual models first. Let us assume that $P_{i,n,f,m}\left(\hat{\Omega}_{m}\right)$ is the probability of person $n$ choosing alternative $i$ in some forecast scenario $f$, according to model $m$, again calculated at the MLEs. Let us further assume that we have worked with model averaging weights at the observation level, i.e. maximising Equation 2 to obtain MLEs for the model weights at the observation level. We can then use these parameters to calculate the model averaging weights for the forecast scenario, i.e. $\pi_{n,f,m}\left(\hat{\Theta}\right)$. The model averaging probability for choosing alternative $i$ in this new setting would then be given by:
$\hat{P}_{i,n,f,MA}=\sum_{m=1}^{M}\hat{\pi}_{n,f,m}\left(\hat{\Theta}\right)P_{i,n,f,m}\left(\hat{\Omega}_{m}\right).$ (5)
2.2 Model averaging for out-of-distribution prediction
Model averaging typically estimates each model on the entire data and then computes the weights by considering how well each model fits for each of the data points, estimating such weights at the person or observation level. This could, for example, include decision-maker characteristics such as age, gender and income and trip characteristics such as purpose and trip distance.
Let us focus on journey distance in particular. Training a model averaging structure where journey distance, say $d_{n,t}$ for person $n$ and trip $t$, is used as a key characteristic in Equation 4 would allow the model to learn how different sub-models are better suited for predicting trips of specific distances within the bounds of the estimation data. However, if the individual sub-models have been estimated on the entire data, then model averaging by definition cannot gain any insights about how the performance might differ across models out-of-distribution, and thus what weights should be given to them outside the range of distances used in estimation. As a result, the weights will simply be an extrapolation based on the estimates for $\gamma_{n,t,m}$ obtained on the training data. For example, imagine that our model would find that the weight assigned to model $m$ increases with distance. If making predictions out-of-distribution, the model would follow the same trend, and for trips with distances below the lower bound of the estimation data, the weight assigned to model $m$ would decrease, while for trips with distances above the upper bound of the estimation data, the weight assigned to model $m$ would increase.
In the present paper, we propose a different approach that enables us to examine the weight assigned to individual models in out-of-distribution predictions. Again, using distance as an example, let us assume that for the available estimation data, we have minimum and maximum distances $d_{min}$ and $d_{max}$, i.e. $D_{data}=\left[d_{min},d_{max}\right]$. We would then define a narrower interval $D_{train}=\left[d_{a},d_{b}\right]$, where $d_{min}<d_{a}<d_{b}<d_{max}$. In the estimation of the individual sub-models, we then use only a subset of the range of the available data, specifically those observations where $d_{n,t}\in D_{train}$. The model averaging structure on the other hand will be estimated on the entire data, i.e. still maximising e.g. Equation 2. This means that the probabilities for the individual models will be calculated for all trips in the data, but using parameters $\hat{\Omega}_{m}$ obtained from optimisation on only a part of the data (where $d_{n,t}\in D_{train}$). Model averaging learns the role of trip characteristics in determining which model should obtain more weight for a specific trip. These are grouped together in the vector $z_{n,t}$ in Equation 4. By splitting the data in the way described above, we can now define new variables to be used in Equation 4 that help characterise whether a given observation is within $D_{train}$, and if not, how far outside the training data it is located.
The above approach directly allows the model averaging procedure to learn how individual sub-models perform outside the range of distances on which they have been trained and how the weight given to each should, as a result, change the further we move away from that interval. This, in turn, then also means that after estimating the model averaging structure, predictions can be made outside the range $D_{data}$, which is of key interest in travel demand forecasting. The model averaging procedure will have learned how the weight assigned to different models changes outside the estimation data and can then apply that knowledge in out-of-distribution prediction. For example, the model averaging process might determine that models with a stronger behavioural foundation suffer a smaller drop in performance outside the estimation data and will thus gain weight in out-of-distribution prediction.
It should be clear that an analyst will need to make important trade-offs in this process. By increasing the gaps $d_{a}-d_{min}$ and $d_{max}-d_{b}$, the analyst will increase the ability of the model averaging structure to learn about out-of-distribution weights. But this will come at the cost of reducing the width of $D_{train}$ and the size of the sample used for training the individual sub-models, reducing the ability to obtain robust estimates of the influence of level-of-service variables, for example.
Hitherto model-averaging studies have primarily relied on logistic regression models, as shown in (4), often with linear specifications. Because of this, such models are ill-equipped to capture complex relations between explanatory variables used for the model averaging and the likelihood of the model $\pi_{n,m}$. However, when it comes to out-of-distribution prediction using model-averaging, nonlinear effects, e.g., of distance, can be expected. Therefore, in this study, we propose using an MLP for the model averaging, which is much more flexible and can learn nonlinearities and interactions between explanatory variables from the data. More specifically, the MLP is trained to learn the function $g()$, which takes the model probabilities, $P_{n,m}$ and explanatory variables $Z_{n}$ to produce the model weight: ${\pi}_{i,n,m}=g(\omega\mid P_{n,m},z_{n})$.
3 Case studies
In this paper, we illustrate the performance of the proposed approach in two case studies. In what follows, we first describe the two datasets, before talking about how the data was divided for the purpose of the present paper. We then talk about the individual sub-models and the specification and estimation of the model averaging structure. We finally look at the results.
3.1 Data
We use two different revealed preference datasets in this study, both focussing on mode choice.
3.1.1 DECISIONS data
The first dataset comes from a large-scale survey conducted as part of the DECISIONS project carried out by the Choice Modelling Centre at the University of Leeds (Calastri et al., 2020). We specifically use the GPS tracking component of this survey, where the data used for this work corresponds to the observed mode choice behaviour.
After extensive data cleaning and data enrichment (Tsoleridis et al., 2022), 12,524 trips made by 540 individuals remained. For each trip, individuals travelled by one of six modes: car (48%), bus (15%), rail (5%), taxi (3%), cycling (3%) or walking (26%). The data covers a very wide range of trip distances, going from just under 100 metres to just under $106$ kilometres, with a mean distance of $8.2$ kilometres and a median distance of $4.5$ kilometres.
The attributes of the alternatives used in the models for the present paper include in-vehicle travel time, out-of-vehicle travel time, and travel cost.
3.1.2 LPMC data
The second dataset is the London mode choice data compiled by Hillel et al. (2018), referred to hereafter as LPMC (London passenger mode choice).
We use a sample of $80,943$ trips made by $31,921$ individuals. This dataset contains four alternatives: walking (18%), cycling (3%), public transport (35%, grouping together bus and rail) and driving (44%, grouping together private car and taxi rides). The data again covers a very wide range of trip distances, going from just under 100 metres to just over $40$ kilometres, with a mean distance of $4.6$ kilometres and a median distance of $2.8$ kilometres.
Attributes of the alternatives used in the models include in-vehicle travel time, out-of-vehicle travel time, interchanges, and travel cost, along with two socio-demographic variables, namely car ownership and driving license status.
3.2 Estimation and validation data
As outlined in the earlier discussions, our approach to model averaging relies on the idea of excluding part of the distance distribution in the training of the individual sub-models, by setting $d_{a}>d_{min}$ and $d_{b}<d_{max}$. For the purpose of testing the performance of the model averaging approach, we go one step further in our empirical work, by excluding very short and very long trips from the training of the model averaging structure as well, thus allowing us to validate the out-of-distribution performance.
Formally, we have that the range of distances in the data is given by $D_{data}=\left[d_{min},d_{max}\right]$. The range used in training the individual sub-models is given by $D_{train,sub}=\left[d_{a,sub},d_{b,sub}\right]$. Finally, the model averaging models are estimated on $D_{train,MA}=\left[d_{a,MA},d_{b,MA}\right]$. Overall, the relationship holds that $d_{min}<d_{a,MA}<d_{a,sub}<d_{b,sub}<d_{b,MA}<d_{max}$. For the purposes of our analysis, we set $d_{a,MA}$ and $d_{b,MA}$ to be at the 10th and 90th percentiles of the sample distance distribution $D_{data}$, while $d_{a,sub}$ and $d_{b,sub}$ are set at the 20th and 80th percentiles of the sample distance distribution $D_{data}$.
In addition to these divisions, we further retain $20\%$ of any trips, independent of their distance, for out-of-sample validation [^1] The splitting of the sample into estimation, validation (and additionally test data for MLP and XGB) was done at the observation rather than person/household level. This might be seen as a limitation, however, none of the models we used explicitly accommodated any individual-specific effects, and information on IDs was not provided to the MLP and XGB training. We thus believe the impact of the way we have split the data to be negligible.. The resulting division of the data is explained in Figure 1, with the resulting sample sizes shown Table 1, where the larger sample validation size for the two extreme segments is a result of this data being used for validation only. In practical work, analysts would be unlikely to not include all deciles (intervals of 10% width) in the MA estimation work and would likely also use a smaller validation sample - we made these decisions with a view to validating the model idea.
Figure 1: Estimation and validation data
Table 1: Estimation and validation data sample sizes for case studies
DECISIONS data LPMC data
Distance (km) Estimation Validation Distance (km) Estimation Validation
Individual MA Individual MA
out-of-distribution for MA (below, i.e. $d_{min}tod_{a,MA}$) <1.33 0 0 1,253 <0.75 0 0 8,092
out-of-distribution for sub-models (below, i.e. $d_{a,MA}$ to $d_{a,sub}$) 1.33 to 1.95 0 1,026 228 0.75 to 1.11 0 6,508 1,580
Within distribution for MA and sub-models 1.95 to 11.95 5,952 1,560 1.11 to 7.41 38,855 9,696
out-of-distribution for sub-models (above, i.e. $d_{b,sub}$ to $d_{b,MA}$) 11.95 to 19.52 0 993 259 7.41 to 11.22 0 6,481 1,623
out-of-distribution for MA (above, i.e. $d_{b,MA}$ to $d_{max}$) 19.52 to 105.9 0 0 1,253 11.22 to 40.1 0 0 8,108
Total 5,952 2,019 4,553 38,855 12,989 29,099
3.3 Model types
The following individual models are used in our analysis, combining structures from choice modelling, mathematical psychology and machine learning. Many other model types could have been considered [^2] A notable omission in our list are mixture models such as mixed logit or latent class - any advantages that such models have in estimation largely/completely disappear in prediction as the random heterogeneity needs to be averaged out. The only exception comes with the use of posterior distributions which are not applicable in the context of out-of-sample prediction. - we limited ourselves to this set with a view to demonstrating the potential benefit of our proposed approach.
Nested Logit (NL):
Our NL models (cf. Daly and Zachary, 1978) use a full set (i.e. $J-1$) of alternative specific constants (ASC). For continuous attributes, we use a linear plus logarithmic specification, i.e. $V=...+\beta_{x}x+\beta_{log-x}log(x)$ to capture non-linearity in sensitivities. For the DECISIONS data, we use three nests, grouping car together with taxi, bus with train, and walking with cycling. For the LPMC data, we use two nests, grouping car together with bus, and walking together with cycling, but this structure ended up collapsing to a Multinomial Logit (MNL) model.
Decision field theory (DFT):
DFT is a dynamic, stochastic model, introduced by Busemeyer and Townsend (1993). The key idea of the DFT model is that the preferences for different alternatives update over time whilst the decision-maker considers the different alternatives and their attributes. We use the implementation of Hancock et al. (2021), with the valence functions following the same specifications as the utility in NL.
MultiLayer Perceptron (MLP):
This model comprises an input layer with input nodes, one or more hidden layers with hidden nodes, and an output layer with output nodes. In this model, signals propagate forward through the links connecting the nodes. The links have numeric weights $w$, which are learned from the data. At each link, the weights are multiplied by the input value from the previous nodes. At the node, the sum of the inputs is taken, and an activation function is applied. The outcome is propagated to the nodes in the next layer. We use tanh activation functions in all hidden layers, except the final one where a linear layer is used. In the output layer, a Softmax function (i.e. a logit) is applied to produce choice probabilities for each alternative.
XGBoost (XGB):
The XGB model comprises a series of sequentially applied decision trees. A decision tree is a sequence of simple IF-THEN rules, optimised to classify data accurately. In the XGB model, each decision tree in the series ‘corrects’ the mispredictions of the models before it. This process is referred to as ‘boosting’. The term ‘gradient’ refers to the fact that each successive tree is fitted using the gradient of the loss function—effectively via gradient descent. Unlike standard Gradient Boosting Machines (GBM), XGBoost includes further optimisations, such as second-order gradient information and advanced regularisation, to improve speed and mitigate overfitting.
The NL and DFT models were coded and estimated using Apollo (cf. Hess and Palma, 2019), while the MLP and XGB models were implemented in Python using the SciKit-learn (Pedregosa et al., 2011) and XGBoost (Chen and Guestrin, 2016) libraries respectively. A hyperparameter search was conducted for both models to optimise their performance.
For the MLP, the hyperparameter search considered the number of hidden layers, the number of nodes per layer, the learning rate, and the L2 regularisation strength. The number of hidden layers and nodes per layer influences the model’s capacity to learn complex patterns, while the learning rate controls how quickly the model updates its weights during training. L2 regularisation helps prevent overfitting by penalising large weight values. Since the best-performing hyperparameters were highly similar across both datasets, we opted to use a single configuration. The best-performing configuration found was two hidden layers with 30 nodes each, a learning rate of 0.001, and an L2 regularisation strength of 0.1. Because each training run of an MLP can yield slightly different outcomes, we trained the model 100 times using the best-performing configuration and averaged their predictions to account for this variability.
For the XGBoost model, the hyperparameter search focused on selecting a subset of features for each tree split, determining the minimum loss reduction needed for further node splitting, setting the maximum tree depth, and specifying the minimum sum of instance weights required in a child node. The first controls the diversity of the model by randomly selecting a fraction of features at each split. The second acts as a pruning mechanism by restricting splits that do not provide sufficient improvement. The third helps balance model complexity and the risk of overfitting. The fourth serves as a regularisation measure, preventing excessively complex splits by enforcing a minimum requirement for node creation. The final parameter determines the fraction of training samples used in each boosting iteration. Similar to the MLPs, since the optimal hyperparameters were nearly identical for both datasets, we opted to use a single configuration. The best-performing values selected were a feature sampling rate of 0.7, a pruning threshold of 2, a maximum tree depth of 2, a minimum child node weight requirement of 1, and a subsampling rate of 1333This is rather shallow, especially in comparison with Hillel (2021), but was the outcome of the hyperperformance tunning on the subset of the data we used.. As with the MLP model, we repeated the training 100 times using the optimal configuration and averaged their predictions to account for variability in the model training. This is common practice, as when training MLPs, a substantial share gets stuck in inferior solutions.
3.4 Model averaging: specification and estimation
Unlike the MLP models implemented for the training on the data sets (see Section 3.3), the MLP for the model averaging is implemented in PyTorch Paszke et al. (2019). This is needed because the loss function (see equation (2)) requires a custom implementation, which is not possible in SciKit-learn. To find the best hyperparameters, hyperparameter tuning was conducted using the following parameters: batch size, number of hidden layers and nodes, learning rate and L2 regularisation. For the DECISIONS data set, the best-performing model has two hidden layers with 60 nodes each, a batch size of 128, a learning rate of 0.005, and an L2 regularisation strength of 0.0001. For the LPMC data set, the best-performing model has two hidden layers with 60 nodes each, a batch size of 64, a learning rate of 0.001, and an L2 regularisation strength of 0.0001.
We repeated the training 100 times using the optimal configuration. As we aim for high prediction performance, we retained only the top 20 percent of models [^4] This selection was based on the log-likelihood. For this, we used the full data, i.e. training and test data, but excluded the validation data., discarding those that became trapped in poor local maxima. This approach, thus, differs slightly from the one used for the training of the data-driven models, where we also repeated training 100 times but averaged predictions across all models to encourage diversity between models.
Figure 2 shows the structure for our model averaging approach. As shown in the figure, the model averaging MLP only takes the trip distance as input to predict the weights given to each sub-model. The model-averaging probability is computed as a weighted sum, combining the predictions for the trip by each sub-model and the weights assigned to each sub-model by the model averaging MLP.
Figure 2: Model averaging structure
3.5 Empirical results
In this section, we present the empirical results for the two case studies. We first focus on the performance of the individual sub-models, comparing this across distance segments and also discussing out-of-sample performance. We then look at the results for the model averaging structure.
3.5.1 Performance of individual sub-models
Table 2 presents the performance of the different models on the estimation sample, covering distance segments (deciles) 3 to 8, as illustrated in Figure 1.
We see that for the DECISIONS data, DFT obtains a higher log-likelihood (LL) than NL, while MNL outperforms DFT in the LPMC data. In both studies, the two machine learning approaches obtain much higher LL than NL/MNL and DFT. Proportionally, the difference is larger in the DECISIONS case study than in the LPMC one, where in the former, MLP also has a bigger advantage over XGB. No formal statistical tests are carried out to compare the different model structures as these are not applicable to the machine learning structures.
Table 2: Final log-likelihood of different model structures (covering distance segments 3-8) of the estimation set
Model DECISIONS LPMC
NL/MNL -2,540.83 -28,782.34
DFT -2,522.93 -28,794.57
MLP -2,095.93 -27,605.21
XGB -2,254.93 -27,635.04
As a next step, Figure 3 looks at the performance of these models across the different distance segments obtained by splitting the data into deciles of the distance distribution. We do this separately for the estimation and validation data. A point to note when studying these results is that for the individual models, only segments 3 to 8 were used in estimating the model parameters, but the probabilities in segments 2 to 9 for the individual models also contributed to the model averaging, as explained in Figure 1. The values shown for segments 2 and 9 in the left two graphs in Figure 3 thus in effect also relate to out-of-distribution validation for the individual models.
Figure 3: Estimation and validation performance
Starting with the graphs on the left, we see that, across the two datasets, the advantage for MLP and XGB that we noted at the overall level in Table 2 apply in each of the 6 distance segments used in estimation (i.e. segments 3-8). Across the distance segments, there are some where XGB outperforms MLP (reversing the overall trend), while the overall trend between DFT and NL/MNL is also reversed in some cases. These findings already show that different models perform differently well across trip distances. Interestingly, while the LPMC results show that mode choice is easier to predict for longer trips (across segments 3-8), no clear trend emerges for the DECISIONS data.
We next turn to segments 2 and 9 in the left two graphs, giving us initial insights into out-of-distribution performance, remembering that the values shown here are the predicted probabilities for the chosen alternative, using the model parameters obtained when estimating the models on segments 3 to 8. In the DECISIONS data, we see all models perform very similarly in segment 2 (except for XGB), while in segment 9, NL clearly outperforms the data-driven models (MLP and XGB) and DFT. In the LPMC models, we see that MLP performs best in both segments 2 and 9, but the differences are very small.
We next turn to the validation results, shown in the right two graphs in Figure 3. We first focus on segments 3-8 with a view to understanding potential overfitting to the estimation data. For the DECISIONS data, we see some evidence of improved out-of-sample (vs within-sample) performance for the NL and DFT, with an average gain in per observation LL of $0.02752$ and $0.01811$, respectively, where this is most notable in segment 8. For MLP and XGB, we see small drops, with changes of $-0.0134$ and $-0.0058$, respectively. In the LPMC models, the changes are smaller, with gains in per observation LL of $0.0098$ for MNL, $0.01$ for DFT, $0.0043$ for MLP, and a drop by $-0.0008$ for XGB. Overall, these results raise no concerns about overfitting to the estimation data, at least within the distribution of distances covered by the estimation data.
The findings for segments 2 and 9 on the validation data largely confirm those we saw in the two graphs on the left in Figure 3. We finally turn to segments 1 and 10. From a behavioural perspective, we see that for the DECISIONS data, mode choice on very short tips is much easier to predict. For short trips in the LPMC dataset, there is a clear advantage for the behavioural models. For the longest trips (segment 10), the trend is less clear. In the DECISIONS data, MLP performs comparatively well. In the LPMC data, both data-driven models do comparatively well, while MNL suffers a big drop.
3.5.2 Model averaging results
In model averaging, data covering segments 2-9 were used, but with model parameters for the individual models estimated on segments 3-8 only (cf. discussion around Figure 1). Table 3 shows the performance (in terms of LL on the estimation set) of the four individual models, alongside that of the model averaging (MA) structure. In both datasets, the MA structure clearly obtains a higher LL than any of the four individual models, although in LPMC, the difference to the best-fitting sub-model (MLP) is only around 0.13%. This in itself is simply a reflection of the overall dominance of MLP over the other three model types in segments 2-9 for this data, and further insights will be gained in model validation. A potential reason for this dominance could be the larger sample size compared to the DECISIONS data, which could enable MLP and XGB to better learn the data generating process.
Table 3: Log-Likelihood for model averaging: individual components and overall structure (covering distance segments 2-9), with gains by MA over individual models shown in brackets
Model DECISIONS LPMC
NL/MNL -3,397.316 (13.62%) -38,383.5 (3.91%)
DFT -3,451.558 (14.98%) -38,366.13 (3.86%)
MLP -3,019.805 (2.83%) -36,931.39 (0.13%)
XGB -3,202.91 (8.38%) -37,228.17 (0.92%)
MA -2,934.45 -36,884.39
The inclusion of segments 2 and 9 allows the model averaging structure to learn about out-of-distribution performance and hence guide the process of linking the weights given to individual models to the difference between the distance for that trip and the data used in estimation (i.e. segments 3 to 8). As explained in Section 3.4, a very flexible neural network was used to estimate this relationship, with no a priori assumptions about the functional form. The resulting relationship is illustrated in Figure 4.
For both datasets, we see that the MA process gives almost the entire weight to MLP and XGB for any trips with a distance contained within $D_{train,sub}$, i.e. distances on which the individual sub-models were estimated. In both datasets, XGB gets higher weight for short trips within that part of the distance distribution compared to long trips, while for longer trips, almost the entire weight goes to MLP. These findings are entirely in line with the pattern of prediction performance obtained for individual models, as shown on the left hand side in Figure 3.
Once we go outside the $D_{train,sub}$ interval, the pattern changes. In both datasets, we observe that for trips with distances below $d_{a,sub}$ (i.e. segments 1&2), the weight for behavioural models increases. In the DECISIONS data, only a small share goes to NL, with quite equal weight for the other three models. On the other hand, in LPMC, MNL essentially obtains all the weight. For longer trips, i.e. above $d_{b,sub}$, NL gets the vast majority of the weight in the DECISIONS data, with the rest going to DFT. For LPMC, MLP retains a large weight, but this decreases while the weight for MNL and DFT increases with distance.
Figure 4: Model averaging: weights given to individual models
The final step in our analysis of results concerns the performance of the model averaging structure on the validation data. Table 4 shows the LL on the validation data, covering the entire range of distances, for the four sub-models and for model averaging. As in estimation, we see that the model averaging structure obtains the least negative LL. What is striking however is the contrast with the results on the estimation data, i.e. Table 2. With the exception of DFT on the LPMC data, the gains by model averaging over the sub-models are larger on the validation data than on the estimation data. This provides initial evidence to support the notion that model averaging is less susceptible to over-fitting on the estimation data than data-driven models.
Table 4: Log-Likelihood for model averaging: individual components and overall structure on validation data (covering distance segments 1-10), with gains by MA over individual models shown in brackets
Model DECISIONS LPMC
NL/MNL -1,948.29 (15.32%) -22,233.56 (14.1%)
DFT -1,980.411 (16.7%) -19,414.46 (1.62%)
MLP -1,722.19 (4.21%) -19,869.84 (3.88%)
XGB -1,895.97 (12.99%) -22,131.08 (13.7%)
MA -1,649.75 -19,099.30
There are two possible influences that could drive the differences between Table 3 and Table 4. The first of these is the potential that model averaging reduces the risk of over-fitting to the estimation data, which can be tested by comparing the performance on segments 2-9 between the estimation and validation data. The second is the potential that model averaging offers improved out-of-distribution performance, which can be tested by comparing the performance on segments 1 and 10 between the sub-models and the model averaging structure.
To support this discussion, Figure 5 shows the per observation log-likelihoods in the different segments for estimation and validation. We first see that, in estimation, model averaging outperforms (or equals) the four sub-models in 4 out of 8 estimation segments for the DECISIONS data, and 7 out of 8 for the LPMC data. When we turn to the validation data, model averaging performs best (or joint best) in 9 out of 10 segments for the DECISIONS data, and 8 out of 10 segments for the LPMC data. These per segment statistics go hand in and hand with the best overall LL obtained by model averaging as reported in Table 3 and Table 4.
Figure 5: MA Estimation and validation performance
In relation to over-fitting to the estimation data, we see that, for all four sub-models as well as for the model averaging structure, there are segments where the per observation LL increases and segments where it decreases. No clear pattern emerges in the comparison across models and segments. Turning to the second potential reason, namely out-of-distribution performance, we observe that for the DECISIONS data, model averaging is outperformed slightly by NL in segment 1, but obtains by far the best performance in segment 10, where its combined performance on segments 1 and 10 is much higher than for any of the other models (-0.3419 vs -0.3628 for MLP). For the LPMC data, a similar story emerges. Model averaging is outperformed slightly by NL in segment 1, and slightly by DFT in segment 10, but overall performs the best of all models when looking at segments 1 and 10 together (-0.6078 vs -0.6153 for DFT).
Overall, these findings lead to the conclusion that there is little difference across the models in terms of out-of-sample performance (when looking at segments 2-9), aside from the potential suggestion that the risk of overfitting is reduced with more behavioural models, as already shown in Figure 3. There is no particular benefit for model averaging over other models in this context (given the dominance of the data-driven models), and nor was this the expectation. On the other hand, there is clear evidence that our model averaging approach offers benefits in terms of out-of-distribution performance, which was its intended aim.
4 Conclusions
This work has taken an important step forward in combining insights from different modelling approaches for travel demand forecasting. Specifically, we have shown that different models predict choices differently well depending on how far away from the estimation sample the prediction takes place. This result is not surprising in itself but is quantified by our work.
The main contribution comes in the development of an MLP-based model averaging approach that estimates weights for different models as a function of the distance away from the estimation data. Across two different case studies looking at mode choice, we see that data-driven machine learning approaches obtain almost the entire weight when making predictions for trips in areas similar to the trips in the estimation data. However, once we move outside that area and look at out-of-distribution prediction, the model averaging approach increases the weight to the models with behavioural underpinnings, sometimes substantially so.
In terms of prediction performance, we observe that model averaging offers the best overall performance on both the estimation and validation data. Crucially, this advantage is at its largest when making out-of-distribution predictions, in our case, predicting mode choice for trips whose distance is outside the range covered by the estimation data.
There are, of course, some opportunities for further refinement.
First, in our implementation, we used the same data to estimate the sub-models and the model averaging. In case of overfitting by one of the sub-models (an issue that did not arise in our case), this could create the risk that model averaging assigns disproportional weight to that model, which would jeopardise the benefits of the approach. In studies with larger sample sizes, a further splitting of the data could be considered.
Second, in applying our proposed model averaging approach, an analyst needs to define what distribution means. In our work, we have focussed on the obvious example of trip distance. A core opportunity for future work is thus the investigation of out-of-distribution performance considering attributes other than trip distance alone, for example, going outside the range of socio-economic characteristics covered by the estimation data, or combinations of socio-economic characteristics not covered in the estimation data. Out-of-distribution could also take on a multivariate nature, thus covering multiple trip characteristics at the same time.
Finally, we used MLP for the model averaging structure. Other meta-models, such as hybrid architectures approaches like TasteNetMNL (cf. Han et al., 2022), are an interesting avenue for future research.
Acknowledgements
Stephane Hess acknowledges the support of the Delft Excellence Fund, as well as the European Research Council through the advanced grant 101020940-SYNERGY. Sander van Cranenburgh acknowledges the support of the TU Delft AI Labs programme. The authors are thankful to Thomas Hancock, Tim Hillel and Georges Sfeir for comments on an earlier draft.

### v2

@@
 Across two case studies, we show that while data-driven approaches excel in predicting mode choice for trips within the distance bands used in estimation, beyond that range, the picture is fuzzy.
-To leverage the relative advantages of the different model families and capitalise on the notion that multiple ‘weak’ models can result in more robust models, we put forward the use of a model averaging approach that allocates weights to different model families as a function of the distance between the characteristics of the trip for which predictions are made, and those used in model estimation.
+To leverage the relative advantages of the different model families and capitalise on the notion that multiple ‘weak’ models can result in more robust models, we put forward the use of a model averaging approach that allocates weights to different model families as a function of the distance between the characteristics of the trip for which predictions are made, and those used in model estimation - in our case, we use trip distance as the sole characteristic for this.
 Overall, we see that the model averaging approach gives larger weight to models with stronger behavioural or econometric underpinnings the more we move outside the interval of trip distances covered in estimation.
-Across both case studies, we show that our model averaging approach obtains improved performance both on the estimation and validation data, and crucially also when predicting mode choices for trips of distances outside the range used in estimation.
+Across both case studies, we show that our model averaging approach obtains improved performance both on the estimation and test data, and crucially also when predicting mode choices for trips of distances outside the range used in estimation.
 Keywords: choice modelling; forecasting; machine learning; mathematical psychology; mode choice; model averaging
@@
 As some of the settings we wish to make predictions for may be outside the range of level-of-service values that the models have been estimated on, it is important for forecasting models to also generalise out-of-distribution.
-Travel behaviour analysts are increasingly interested in a more diverse set of models, covering econometric structures, models from mathematical psychology, and machine learning techniques.
+Travel behaviour analysts are increasingly interested in a more diverse set of models, covering not just econometric structures, but also models from mathematical psychology (e.g.
+Hancock et al., 2021, 2020a), and increasingly machine learning techniques (e.g.
+Alwosheel et al., 2021; Wong and Farooq, 2021; Sfeir et al., 2022; Yan et al., 2025; Wang et al., 2020).
+This is especially true given the growing reliance on highly granualar data on travel behaviour, for example that collected using GPS tracking (e.g.
+Liu et al., 2024; Tsoleridis et al., 2025).
 With the typical model evaluation approach focussing on the ability of models to replicate the behaviour in the empirical data as well as possible, it should then come as no surprise that researchers in the travel behaviour field are increasingly attracted by the comparatively good prediction performance of machine learning approaches (cf.
@@
 Second, we aim to develop an approach that improves prediction performance in forecasts through a model averaging approach that assigns different weights to different models depending on this “distance".
-In our work, we use trip distance as the key variable, but the approach works with any variable that goes out of distribution, not only physical distance.
+In our work, we use trip distance as the key variable, but the approach works with any variable that goes out of distribution, not only geographic distance.
 We employ a neural network for the model averaging.
@@
 In model averaging (see e.g.
-Hancock et al., 2020), we first estimate $M$ different models on the data, with model $m$ giving a likelihood $L_{n,t,m}\left(\Omega_{m}\right)$ for the choice observed by decision maker $n$ in choice situation $t$, using a vector of estimated model parameters $\Omega_{m}$.
+Hancock et al., 2020b), which can be viewed as a form of stacking, first introduced in the seminal work of Wolpert (Wolpert, 1992), which established the conceptual basis for ensemble methods combining multiple predictive models.
+Specifically, we first estimate $M$ different models on the data, with model $m$ giving a likelihood $L_{n,t,m}\left(\Omega_{m}\right)$ for the choice observed by decision maker $n$ in choice situation $t$, using a vector of estimated model parameters $\Omega_{m}$.
 Let us further define $L_{n,m}\left(\Omega_{m}\right)$ to be the likelihood for the entire sequence of $T_{n}$ choices observed for person $n$ according to model $m$.
@@
 In maximising the model averaging likelihood, an analyst thus, in fact, does not require the model parameters for the individual models, only the probabilities of each observed choice.
-In applying the model to predict choices out-of-sample, an analyst then again requires the predicted choice probabilities from the individual models first.
-Let us assume that $P_{i,n,f,m}\left(\hat{\Omega}_{m}\right)$ is the probability of person $n$ choosing alternative $i$ in some forecast scenario $f$, according to model $m$, again calculated at the MLEs.
-Let us further assume that we have worked with model averaging weights at the observation level, i.e.
-maximising Equation 2 to obtain MLEs for the model weights at the observation level.
-We can then use these parameters to calculate the model averaging weights for the forecast scenario, i.e.
-$\pi_{n,f,m}\left(\hat{\Theta}\right)$.
-The model averaging probability for choosing alternative $i$ in this new setting would then be given by:
+[…]
 $\hat{P}_{i,n,f,MA}=\sum_{m=1}^{M}\hat{\pi}_{n,f,m}\left(\hat{\Theta}\right)P_{i,n,f,m}\left(\hat{\Omega}_{m}\right).$ (5)
@@
 This could, for example, include decision-maker characteristics such as age, gender and income and trip characteristics such as purpose and trip distance.
-Let us focus on journey distance in particular.
+In the present paper, we focus on trip distance in particular.
+This is to some extent a pragmatic choice given the proof-of-concept nature of our work.
+It allows us to very easily split the data and define what constitutes within-distribution and out-of-distribution, with a view to demonstrating the performance of the model averaging approach.
+In addition however, trip distance has a number of appealing characteristics in this context.
+First, distance is not typically used as an explanatory variable in the mode choice models themselves, which instead rely on the closely correlated variables such as time and cost.
+Second, distance can explain much of the similarity between trips, along multiple dimensions at the same time.
+This way, the model averaging structure can learn which model works well for what trip distance, without having to rely on multiple covariates, given that this single variable condenses other information into a single factor.
 Training a model averaging structure where journey distance, say $d_{n,t}$ for person $n$ and trip $t$, is used as a key characteristic in Equation 4 would allow the model to learn how different sub-models are better suited for predicting trips of specific distances within the bounds of the estimation data.
@@
 We specifically use the GPS tracking component of this survey, where the data used for this work corresponds to the observed mode choice behaviour.
-After extensive data cleaning and data enrichment (Tsoleridis et al., 2022), 12,524 trips made by 540 individuals remained.
-For each trip, individuals travelled by one of six modes: car (48%), bus (15%), rail (5%), taxi (3%), cycling (3%) or walking (26%).
+After extensive data cleaning and data enrichment (Tsoleridis et al., 2022), 10,990 trips made by 415 individuals remained.
+For each trip, individuals travelled by one of six modes: car (48%), bus (14%), rail (5%), taxi (3%), cycling (4%) or walking (26%).
 The data covers a very wide range of trip distances, going from just under 100 metres to just under $106$ kilometres, with a mean distance of $8.2$ kilometres and a median distance of $4.5$ kilometres.
@@
 Attributes of the alternatives used in the models include in-vehicle travel time, out-of-vehicle travel time, interchanges, and travel cost, along with two socio-demographic variables, namely car ownership and driving license status.
-3.2 Estimation and validation data
+3.2 Estimation and test data
 As outlined in the earlier discussions, our approach to model averaging relies on the idea of excluding part of the distance distribution in the training of the individual sub-models, by setting $d_{a}>d_{min}$ and $d_{b}<d_{max}$.
@@
 For the purposes of our analysis, we set $d_{a,MA}$ and $d_{b,MA}$ to be at the 10th and 90th percentiles of the sample distance distribution $D_{data}$, while $d_{a,sub}$ and $d_{b,sub}$ are set at the 20th and 80th percentiles of the sample distance distribution $D_{data}$.
-In addition to these divisions, we further retain $20\%$ of any trips, independent of their distance, for out-of-sample validation [^1] The splitting of the sample into estimation, validation (and additionally test data for MLP and XGB) was done at the observation rather than person/household level.
-This might be seen as a limitation, however, none of the models we used explicitly accommodated any individual-specific effects, and information on IDs was not provided to the MLP and XGB training.
-We thus believe the impact of the way we have split the data to be negligible..
-The resulting division of the data is explained in Figure 1, with the resulting sample sizes shown Table 1, where the larger sample validation size for the two extreme segments is a result of this data being used for validation only.
-In practical work, analysts would be unlikely to not include all deciles (intervals of 10% width) in the MA estimation work and would likely also use a smaller validation sample - we made these decisions with a view to validating the model idea.
-Figure 1: Estimation and validation data
-Table 1: Estimation and validation data sample sizes for case studies
+In addition to these divisions, we further retain $20\%$ of any trips, independent of their distance, for out-of-sample test.
+The splitting of the sample into estimation, test (and additionally validation data for MLP and XGB) was done at the person/household level, so as to avoid data leakage.
+In other words, each person only contributed data to either the estimate data or the test data.
+For those individuals who contributed to the estimation data, their observations in distance segments 1 and 10 were thus not used in our analysis.
+The resulting division of the data is explained in Figure 1, with the resulting sample sizes shown Table 1.
+In practical work, analysts would be unlikely to not include all deciles (intervals of 10% width) in the MA estimation work - we made these decisions with a view to validating the model idea.
+Similarly, a real-world application would not set aside the test data, and thus also avoid deleting observations in the estimation data for trips that are in segments 1 and 10 (see the earlier point about data leakage).
+Figure 1: Estimation and test data
+Table 1: Estimation and test data sample sizes for case studies
 DECISIONS data LPMC data
-Distance (km) Estimation Validation Distance (km) Estimation Validation
+Distance (km) Estimation test Distance (km) Estimation test
 Individual MA Individual MA
-out-of-distribution for MA (below, i.e. $d_{min}tod_{a,MA}$) <1.33 0 0 1,253 <0.75 0 0 8,092
-out-of-distribution for sub-models (below, i.e. $d_{a,MA}$ to $d_{a,sub}$) 1.33 to 1.95 0 1,026 228 0.75 to 1.11 0 6,508 1,580
-Within distribution for MA and sub-models 1.95 to 11.95 5,952 1,560 1.11 to 7.41 38,855 9,696
-out-of-distribution for sub-models (above, i.e. $d_{b,sub}$ to $d_{b,MA}$) 11.95 to 19.52 0 993 259 7.41 to 11.22 0 6,481 1,623
-out-of-distribution for MA (above, i.e. $d_{b,MA}$ to $d_{max}$) 19.52 to 105.9 0 0 1,253 11.22 to 40.1 0 0 8,108
-Total 5,952 2,019 4,553 38,855 12,989 29,099
+out of distribution for MA (below, i.e. $d_{min}-d_{a,MA}$ <1.33 0 0 223 <0.75 0 0 1,581
+out-of-distribution for sub-models (below, i.e. $d_{a,MA}$ to $d_{a,sub}$) 1.33 to 1.95 0 836 261 0.75 to 1.11 0 6,446 1,642
+within distribution for MA and sub-models 1.95 to 11.95 5,272 1,322 1.11 to 7.41 38,761 9,790
+out-of-distribution for sub-models (above, i.e. $d_{b,sub}$ to $d_{b,MA}$) 11.95 to 19.52 0 946 153 7.41 to 11.22 0 6,458 1,646
+out of distribution for MA (above, i.e. $d_{b,MA}tod_{max}$ 19.52 to 105.9 0 0 154 11.22 to 40.1 0 0 1,630
+total 5,272 1,782 2,113 38,761 12,904 16,289
 3.3 Model types
 The following individual models are used in our analysis, combining structures from choice modelling, mathematical psychology and machine learning.
-Many other model types could have been considered [^2] A notable omission in our list are mixture models such as mixed logit or latent class - any advantages that such models have in estimation largely/completely disappear in prediction as the random heterogeneity needs to be averaged out.
-The only exception comes with the use of posterior distributions which are not applicable in the context of out-of-sample prediction.
+Many other model types could have been considered [^1] A notable omission in our list is mixture models such as mixed logit or latent class - any advantages that such models have in estimation largely/completely disappear in prediction as the random heterogeneity needs to be averaged out.
+The only exception comes with the use of posterior distributions, which are not applicable in the context of out-of-sample prediction.
 - we limited ourselves to this set with a view to demonstrating the potential benefit of our proposed approach.
@@
 Daly and Zachary, 1978) use a full set (i.e.
-$J-1$) of alternative specific constants (ASC).
+J1J-1) of alternative specific constants (ASC).
 For continuous attributes, we use a linear plus logarithmic specification, i.e.
@@
 The outcome is propagated to the nodes in the next layer.
-We use tanh activation functions in all hidden layers, except the final one where a linear layer is used.
+We use tanh activation functions in all hidden layers, except the final one, where a linear layer is used.
 In the output layer, a Softmax function (i.e.
 a logit) is applied to produce choice probabilities for each alternative.
+To ensure consistency with the behavioural models and enhance predictive performance, availability conditions are explicitly incorporated into the network (thus, treated in the same way as in traditional choice models).
 XGBoost (XGB):
@@
 Unlike standard Gradient Boosting Machines (GBM), XGBoost includes further optimisations, such as second-order gradient information and advanced regularisation, to improve speed and mitigate overfitting.
+To ensure consistency with the behavioural models and to better handle stronger regularisation settings, availability conditions are enforced via base_margin: we set margins to 0 for available alternatives and a large negative value for unavailable ones, ensuring that the model assigns zero probability to non-available alternatives.
+The model is implemented directly with the xgboost package.
 The NL and DFT models were coded and estimated using Apollo (cf.
-Hess and Palma, 2019), while the MLP and XGB models were implemented in Python using the SciKit-learn (Pedregosa et al., 2011) and XGBoost (Chen and Guestrin, 2016) libraries respectively.
-A hyperparameter search was conducted for both models to optimise their performance.
-For the MLP, the hyperparameter search considered the number of hidden layers, the number of nodes per layer, the learning rate, and the L2 regularisation strength.
-The number of hidden layers and nodes per layer influences the model’s capacity to learn complex patterns, while the learning rate controls how quickly the model updates its weights during training.
-L2 regularisation helps prevent overfitting by penalising large weight values.
+Hess and Palma, 2019), while the MLP and XGB models were implemented in Python.
+The MLP is implemented in PyTorch (Pedregosa et al., 2011) using a custom to handle the availability conditions, while the XGBoost model is implemented directly with the xgboost package (Chen and Guestrin, 2016).
+For both models, a hyperparameter grid search was conducted to optimise performance.
+We used a 5-fold Grouped cross-validation procedure to account for the panel structure of the data (see Table 2).
+The hyperparameter ranges for the XGB model were informed by the results of Hillel (2021).
+The number of hidden layers and nodes per layer influences the model’s capacity to learn complex patterns, while the batch size affects the stability and efficiency of the gradient updates.
+The learning rate controls how quickly the model updates its weights during training, while the L2 regularisation helps prevent overfitting by penalising large weight values.
 Since the best-performing hyperparameters were highly similar across both datasets, we opted to use a single configuration.
-The best-performing configuration found was two hidden layers with 30 nodes each, a learning rate of 0.001, and an L2 regularisation strength of 0.1.
-Because each training run of an MLP can yield slightly different outcomes, we trained the model 100 times using the best-performing configuration and averaged their predictions to account for this variability.
-For the XGBoost model, the hyperparameter search focused on selecting a subset of features for each tree split, determining the minimum loss reduction needed for further node splitting, setting the maximum tree depth, and specifying the minimum sum of instance weights required in a child node.
-The first controls the diversity of the model by randomly selecting a fraction of features at each split.
-The second acts as a pruning mechanism by restricting splits that do not provide sufficient improvement.
-The third helps balance model complexity and the risk of overfitting.
-The fourth serves as a regularisation measure, preventing excessively complex splits by enforcing a minimum requirement for node creation.
-The final parameter determines the fraction of training samples used in each boosting iteration.
-Similar to the MLPs, since the optimal hyperparameters were nearly identical for both datasets, we opted to use a single configuration.
-The best-performing values selected were a feature sampling rate of 0.7, a pruning threshold of 2, a maximum tree depth of 2, a minimum child node weight requirement of 1, and a subsampling rate of 1333This is rather shallow, especially in comparison with Hillel (2021), but was the outcome of the hyperperformance tunning on the subset of the data we used..
+The best-performing configuration found was two hidden layers with 30 nodes each, a learning rate of 0.001, a batch size of 250 and an L2 regularisation strength of 0.
+Because each training run of an MLP can yield slightly different outcomes, we trained the model 100 times using the best-performing configuration and averaged its predictions to account for this variability.
+For the XGBoost model, the hyperparameter search included the maximum tree depth, the number of estimators, the minimum loss reduction, the minimum child weight, the subsample fraction, the fraction of features considered per split, regularisation strengths (L1 and L2), and the maximum delta step.
+The maximum tree depth controls the complexity of each tree, with deeper trees allowing more detailed partitions at the risk of overfitting.
+The number of estimators determines how many boosting rounds are performed, thereby influencing the overall capacity of the model.
+The minimum loss reduction acts as a pruning mechanism by preventing splits that do not provide sufficient improvement to the objective function, while the minimum child weight enforces a lower bound on the sum of instance weights required in a leaf, discouraging overly specific partitions.
+The subsample fraction regulates the proportion of training samples used for each boosting round, and the feature subsampling parameter (colsample by tree) increases model robustness by introducing randomness at the feature level.
+L1 and L2 regularisation terms penalise overly large weights and help stabilise the model under strong regularisation.
+Finally, the maximum delta step restricts the weight update size, improving convergence stability in cases of unbalanced data.
+Unlike the MLPs, the optimal hyperparameters for the XGBoost model were not identical across the datasets.
+Both the London and Decision data shared a fixed learning rate of 0.01, a minimum loss reduction of 0.6, a minimum child weight of 20, and a subsample fraction of 0.7.
+However, the London configuration used a feature subsampling rate of 0.75, a maximum tree depth of 6, 1,200 boosting rounds, and regularisation strengths of $L_{1}=20$ and $L_{2}=40$, with a maximum delta step of 6.
+In contrast, the Decision configuration relied on full feature sampling (1.0), a deeper maximum tree depth of 9, 800 boosting rounds, and lighter $L_{1}$ regularisation of 10 (with $L_{2}=40$), combined with a smaller maximum delta step of 4.
 As with the MLP model, we repeated the training 100 times using the optimal configuration and averaged their predictions to account for variability in the model training.
 This is common practice, as when training MLPs, a substantial share gets stuck in inferior solutions.
+Table 2: Hyperparameter spaces considered for the MLP and XGBoost models.
+MLP XGBoost
+Hidden layers: $\{(10,5),(10,10),$ (20,20),(30,30)}\;(20,20),(30,30)\} Batch size: $\{250,500,1000\}$ Learning rate: $\{0.01,0.001,0.0001\}$ $L_{2}$ regularisation: $\{0,0.1,0.5,1\}$ Max.
+tree depth: $\{6,9\}$ Number of estimators: $\{800,1200\}$ Min.
+loss reduction: $\{0.2,0.6\}$ Min.
+child weight: $\{10,20\}$ Subsample fraction: $\{0.7,0.8\}$ Colsample by tree: $\{0.75,1.0\}$ $L_{1}$ regularisation: $\{10,20,40\}$ $L_{2}$ regularisation: $\{10,20,40\}$ Max.
+delta step: $\{4,6\}$
 3.4 Model averaging: specification and estimation
-Unlike the MLP models implemented for the training on the data sets (see Section 3.3), the MLP for the model averaging is implemented in PyTorch Paszke et al.
-(2019).
-This is needed because the loss function (see equation (2)) requires a custom implementation, which is not possible in SciKit-learn.
-To find the best hyperparameters, hyperparameter tuning was conducted using the following parameters: batch size, number of hidden layers and nodes, learning rate and L2 regularisation.
-For the DECISIONS data set, the best-performing model has two hidden layers with 60 nodes each, a batch size of 128, a learning rate of 0.005, and an L2 regularisation strength of 0.0001.
+Similar to the MLP models implemented for the training on the data sets (see Section 3.3), we implemented the MLP for the model averaging in PyTorch Paszke et al.
+(2019) too.
+This is needed because the loss function (see equation (2)) requires a custom implementation, which is not possible in e.g.
+SciKit-learn.
+As explanatory features, we included three distance-based variables: the raw trip distance, its logarithmic transformation, and its squared term.
+While in theory the linear distance alone should suffice, we found that adding the log- and squared transformations provided a small but consistent improvement in model performance.
+To tune the hyperparameters, we performed a grid search using a simple 75%–25% train–test split.
+Ideally, a dedicated validation set would have been used for this purpose—separate from the data employed in training the individual models, see Hastie (2009), who argues that not doing so might result in the meta-model assigning unfairly high weights to models with higher complexity.
+However, we deliberately refrained from further partitioning the data to avoid diminishing the already limited estimation sample and to ensure that the test set remained completely untouched for the final performance evaluation.
+The following parameter spaces are used: batch size (64, 128, 256), number of hidden layers and nodes (two or three hidden layers with each 10, 20, 40, or 60 nodes), learning rate (0.001, 0.005, 0.01) and L2 regularisation (0, 0.0001, 0.001).
+For the DECISIONS data set, the best-performing model has two hidden layers with 20 nodes each, a batch size of 128, a learning rate of 0.005, and an L2 regularisation strength of 0.001.
 For the LPMC data set, the best-performing model has two hidden layers with 60 nodes each, a batch size of 64, a learning rate of 0.001, and an L2 regularisation strength of 0.0001.
@@
 For this, we used the full data, i.e.
-training and test data, but excluded the validation data., discarding those that became trapped in poor local maxima.
+training and validation data, but excluded the test data., discarding those that became trapped in poor local maxima.
 This approach, thus, differs slightly from the one used for the training of the data-driven models, where we also repeated training 100 times but averaged predictions across all models to encourage diversity between models.
@@
 As shown in the figure, the model averaging MLP only takes the trip distance as input to predict the weights given to each sub-model.
-The model-averaging probability is computed as a weighted sum, combining the predictions for the trip by each sub-model and the weights assigned to each sub-model by the model averaging MLP.
+The model-averaging probability is computed as a weighted sum, combining the predictions for the trip by each sub-model and the weights assigned to each sub-model by the model-averaging MLP.
 Figure 2: Model averaging structure
@@
 Table 3 presents the performance of the different models on the estimation sample, covering distance segments (deciles) 3 to 8, as illustrated in Figure 1.
-We see that for the DECISIONS data, DFT obtains a higher log-likelihood (LL) than NL, while MNL outperforms DFT in the LPMC data.
+We see that for the DECISIONS data, NL obtains a higher log-likelihood (LL) than DFT, while DFT outperforms MNL in the LPMC data.
 In both studies, the two machine learning approaches obtain much higher LL than NL/MNL and DFT.
-Proportionally, the difference is larger in the DECISIONS case study than in the LPMC one, where in the former, MLP also has a bigger advantage over XGB.
+Proportionally, the difference is larger in the DECISIONS case study than in the LPMC one.
+In the former, MLP outperforms XGB, while the reverse applies for the latter.
 No formal statistical tests are carried out to compare the different model structures as these are not applicable to the machine learning structures.
@@
 Model DECISIONS LPMC
-NL/MNL -2,540.83 -28,782.34
-DFT -2,522.93 -28,794.57
-MLP -2,095.93 -27,605.21
-XGB -2,254.93 -27,635.04
+NL/MNL -1,809.47 -23,067.21
+DFT -1,816.32 -23,032.42
+MLP -1,614.22 -22,154.49
+XGB -1,650.34 -21,628.70
 As a next step, Figure 3 looks at the performance of these models across the different distance segments obtained by splitting the data into deciles of the distance distribution.
-We do this separately for the estimation and validation data.
+We do this separately for the estimation and test data.
 A point to note when studying these results is that for the individual models, only segments 3 to 8 were used in estimating the model parameters, but the probabilities in segments 2 to 9 for the individual models also contributed to the model averaging, as explained in Figure 1.
 The values shown for segments 2 and 9 in the left two graphs in Figure 3 thus in effect also relate to out-of-distribution validation for the individual models.
-Figure 3: Estimation and validation performance
+Figure 3: Estimation and test performance
 Starting with the graphs on the left, we see that, across the two datasets, the advantage for MLP and XGB that we noted at the overall level in Table 3 apply in each of the 6 distance segments used in estimation (i.e.
 segments 3-8).
-Across the distance segments, there are some where XGB outperforms MLP (reversing the overall trend), while the overall trend between DFT and NL/MNL is also reversed in some cases.
+Across the distance segments, there are some where XGB outperforms MLP in the DECISIONS data, or MLP outperforms XGB in LPMC (reversing the overall trend), while the overall trend between DFT and NL/MNL is also reversed in some cases.
 These findings already show that different models perform differently well across trip distances.
@@
 We next turn to segments 2 and 9 in the left two graphs, giving us initial insights into out-of-distribution performance, remembering that the values shown here are the predicted probabilities for the chosen alternative, using the model parameters obtained when estimating the models on segments 3 to 8.
-In the DECISIONS data, we see all models perform very similarly in segment 2 (except for XGB), while in segment 9, NL clearly outperforms the data-driven models (MLP and XGB) and DFT.
-In the LPMC models, we see that MLP performs best in both segments 2 and 9, but the differences are very small.
-We next turn to the validation results, shown in the right two graphs in Figure 3.
+In the DECISIONS data, we see all models perform very similarly in segment 2, while in segment 9, DFT outperforms XGB.
+In the LPMC models, NL and DFT perform best in segment 2, while MLP does best in segment 9, but differences are again small.
+We next turn to the test results, shown in the right two graphs in Figure 3.
 We first focus on segments 3-8 with a view to understanding potential overfitting to the estimation data.
-For the DECISIONS data, we see some evidence of improved out-of-sample (vs within-sample) performance for the NL and DFT, with an average gain in per observation LL of $0.02752$ and $0.01811$, respectively, where this is most notable in segment 8.
-For MLP and XGB, we see small drops, with changes of $-0.0134$ and $-0.0058$, respectively.
-In the LPMC models, the changes are smaller, with gains in per observation LL of $0.0098$ for MNL, $0.01$ for DFT, $0.0043$ for MLP, and a drop by $-0.0008$ for XGB.
-Overall, these results raise no concerns about overfitting to the estimation data, at least within the distribution of distances covered by the estimation data.
-The findings for segments 2 and 9 on the validation data largely confirm those we saw in the two graphs on the left in Figure 3.
+For the DECISIONS data, we see some evidence of improved out-of-sample (vs within-sample) performance for all models, with an average gain in per observation LL of $0.096$ for NL, $0.095$ for DFT, $0.062$ for MLP, and $0.054$ for XGB.
+The same happens for LPMC models, the same is observed, with improvements by $0.12$ for NL and DFT, $0.11$ for MLP and $0.094$ for XGB.
+Overall, these results do not suggest any overfitting, at least within the distribution of distances covered by the estimation data.
+The findings for segments 2 and 9 on the test data largely confirm those we saw in the two graphs on the left in Figure 3, except for MLP and XGB performing more poorly in the test data for segment 9.
 We finally turn to segments 1 and 10.
@@
 Table 4 shows the performance (in terms of LL on the estimation set) of the four individual models, alongside that of the model averaging (MA) structure.
-In both datasets, the MA structure clearly obtains a higher LL than any of the four individual models, although in LPMC, the difference to the best-fitting sub-model (MLP) is only around 0.13%.
-This in itself is simply a reflection of the overall dominance of MLP over the other three model types in segments 2-9 for this data, and further insights will be gained in model validation.
-A potential reason for this dominance could be the larger sample size compared to the DECISIONS data, which could enable MLP and XGB to better learn the data generating process.
+In both datasets, the MA structure clearly obtains a higher LL than any of the four individual models.
+It is noteworthy that this finding is robust even to variations in the performance of the individual models.
+In other words, the MA structure attains a similarly high log-likelihood even when one or more of the constituent models perform less well.
+This echoes the well-known principle in machine learning that an ensemble of weak learners can together form a strong learner.
+From a practical perspective, this suggests that analysts need not exhaustively optimise each individual model when developing a model-averaging framework.
 Table 4: Log-Likelihood for model averaging: individual components and overall structure (covering distance segments 2-9), with gains by MA over individual models shown in brackets
 Model DECISIONS LPMC
-NL/MNL -3,397.316 (13.62%) -38,383.5 (3.91%)
-DFT -3,451.558 (14.98%) -38,366.13 (3.86%)
-MLP -3,019.805 (2.83%) -36,931.39 (0.13%)
-XGB -3,202.91 (8.38%) -37,228.17 (0.92%)
-MA -2,934.45 -36,884.39
+NL/MNL -3,134.33 (11.15%) -40,210.60 (4.49%)
+DFT -3,061.83 (9.04%) -40,186.25 (4.43%)
+MLP -2,846.66 (2.17%) -38,966.05 (1.44%)
+XGB -2,999.89 (7.16%) -38,882.09 (1.23%)
+MA -2,784.99 -38,404.69
 The inclusion of segments 2 and 9 allows the model averaging structure to learn about out-of-distribution performance and hence guide the process of linking the weights given to individual models to the difference between the distance for that trip and the data used in estimation (i.e.
 segments 3 to 8).
-As explained in Section 3.4, a very flexible neural network was used to estimate this relationship, with no a priori assumptions about the functional form.
+As explained in Section 3.4, a flexible neural network is used to estimate this relationship, with no a priori assumptions about the functional form.
 The resulting relationship is illustrated in Figure 4.
@@
 distances on which the individual sub-models were estimated.
-In both datasets, XGB gets higher weight for short trips within that part of the distance distribution compared to long trips, while for longer trips, almost the entire weight goes to MLP.
-These findings are entirely in line with the pattern of prediction performance obtained for individual models, as shown on the left hand side in Figure 3.
+In both datasets, XGB gets higher weight for short trips within that part of the distance distribution compared to long trips, while for longer trips, almost the entire weight goes to MLP for DECISIONS, while it also increases for LPMC.
+These findings are in line with the pattern of prediction performance obtained for individual models, as shown on the left hand side in Figure 3.
 Once we go outside the $D_{train,sub}$ interval, the pattern changes.
@@
 Figure 4: Model averaging: weights given to individual models
-The final step in our analysis of results concerns the performance of the model averaging structure on the validation data.
-Table 4 shows the LL on the validation data, covering the entire range of distances, for the four sub-models and for model averaging.
-As in estimation, we see that the model averaging structure obtains the least negative LL.
-What is striking however is the contrast with the results on the estimation data, i.e.
+The final step in our analysis of results concerns the performance of the model averaging structure on the test data.
+Table 5 shows the LL on the test data, covering the entire range of distances, for the four sub-models and for model averaging.
+As in estimation, we see that the model averaging structure obtains the highest LL.
+What is notable is the contrast with the results on the estimation data, i.e.
 Table 3.
-With the exception of DFT on the LPMC data, the gains by model averaging over the sub-models are larger on the validation data than on the estimation data.
-This provides initial evidence to support the notion that model averaging is less susceptible to over-fitting on the estimation data than data-driven models.
-Table 4: Log-Likelihood for model averaging: individual components and overall structure on validation data (covering distance segments 1-10), with gains by MA over individual models shown in brackets
+For both datasets, the gains by model averaging over the best performing sub-model are larger on the test data than on the estimation data.
+Table 5: Log-Likelihood for model averaging: individual components and overall structure on test data (covering distance segments 1-10), with gains by MA over individual models shown in brackets
 Model DECISIONS LPMC
-NL/MNL -1,948.29 (15.32%) -22,233.56 (14.1%)
-DFT -1,980.411 (16.7%) -19,414.46 (1.62%)
-MLP -1,722.19 (4.21%) -19,869.84 (3.88%)
-XGB -1,895.97 (12.99%) -22,131.08 (13.7%)
-MA -1,649.75 -19,099.30
+NL/MNL -825.88 (6.12%) -11,764.56 (5.30%)
+DFT -805.78 (3.78%) -11,426.72 (2.50%)
+MLP -804.91 (3.68%) -11,448.70 (2.68%)
+XGB -829.64 (6.55%) -12,082.56 (7.79%)
+MA -775.32 -11,141.55
 There are two possible influences that could drive the differences between Table 4 and Table 5.
-The first of these is the potential that model averaging reduces the risk of over-fitting to the estimation data, which can be tested by comparing the performance on segments 2-9 between the estimation and validation data.
+The first of these is the potential that model averaging reduces the risk of over-fitting to the estimation data, which can be tested by comparing the performance on segments 2-9 between the estimation and test data.
 The second is the potential that model averaging offers improved out-of-distribution performance, which can be tested by comparing the performance on segments 1 and 10 between the sub-models and the model averaging structure.
-To support this discussion, Figure 5 shows the per observation log-likelihoods in the different segments for estimation and validation.
-We first see that, in estimation, model averaging outperforms (or equals) the four sub-models in 4 out of 8 estimation segments for the DECISIONS data, and 7 out of 8 for the LPMC data.
-When we turn to the validation data, model averaging performs best (or joint best) in 9 out of 10 segments for the DECISIONS data, and 8 out of 10 segments for the LPMC data.
-These per segment statistics go hand in and hand with the best overall LL obtained by model averaging as reported in Table 3 and Table 4.
-Figure 5: MA Estimation and validation performance
-In relation to over-fitting to the estimation data, we see that, for all four sub-models as well as for the model averaging structure, there are segments where the per observation LL increases and segments where it decreases.
+To support this discussion, Figure 5 shows the per observation log-likelihoods in the different segments for estimation and test.
+We first see that, in estimation, model averaging outperforms (or equals) the four sub-models in 4 out of 8 estimation segments for the DECISIONS data, and 6 out of 8 for the LPMC data.
+When we turn to the test data, model averaging performs best (or joint best) in 2 out of 10 segments for the DECISIONS data, and 5 out of 10 segments for the LPMC data.
+These per segment statistics go hand in hand with the best overall LL obtained by model averaging as reported in Table 4 and Table 5.
+Figure 5: MA Estimation and test performance
+We earlier raised two possible reasons for why the advantages of MA might be larger on the test data than on the validation data.
+In relation to over-fitting to the estimation data, we see that, for all four sub-models as well as for the model averaging structure, there are segments (out of 2-9) where the per observation LL increases and segments where it decreases.
 No clear pattern emerges in the comparison across models and segments.
-Turning to the second potential reason, namely out-of-distribution performance, we observe that for the DECISIONS data, model averaging is outperformed slightly by NL in segment 1, but obtains by far the best performance in segment 10, where its combined performance on segments 1 and 10 is much higher than for any of the other models (-0.3419 vs -0.3628 for MLP).
+Turning to the second potential reason, namely out-of-distribution performance, we observe that for the DECISIONS data, model averaging is outperformed by NL in segment 1, but obtains by far the best performance in segment 10, where its combined performance on segments 1 and 10 is much higher than for any of the other models (-0.2582 vs -0.2944 for XGB).
 For the LPMC data, a similar story emerges.
-Model averaging is outperformed slightly by NL in segment 1, and slightly by DFT in segment 10, but overall performs the best of all models when looking at segments 1 and 10 together (-0.6078 vs -0.6153 for DFT).
-Overall, these findings lead to the conclusion that there is little difference across the models in terms of out-of-sample performance (when looking at segments 2-9), aside from the potential suggestion that the risk of overfitting is reduced with more behavioural models, as already shown in Figure 3.
-There is no particular benefit for model averaging over other models in this context (given the dominance of the data-driven models), and nor was this the expectation.
+Model averaging is outperformed slightly by NL in segment 1, but obtains by far the best performance in segment 10.
+Again, overall, MA performs the best of all models when looking at segments 1 and 10 together (-0.5793 vs -0.6025 for DFT).
+Overall, these findings lead to the conclusion that there is little difference across the models in terms of out-of-sample performance when looking at segments 2-9.
 On the other hand, there is clear evidence that our model averaging approach offers benefits in terms of out-of-distribution performance, which was its intended aim.
@@
 This work has taken an important step forward in combining insights from different modelling approaches for travel demand forecasting.
-Specifically, we have shown that different models predict choices differently well depending on how far away from the estimation sample the prediction takes place.
+Specifically, we have shown that different models predict choices differently well whether we are within the ranges of the estimation data, or outside.
 This result is not surprising in itself but is quantified by our work.
@@
 Across two different case studies looking at mode choice, we see that data-driven machine learning approaches obtain almost the entire weight when making predictions for trips in areas similar to the trips in the estimation data.
-However, once we move outside that area and look at out-of-distribution prediction, the model averaging approach increases the weight to the models with behavioural underpinnings, sometimes substantially so.
-In terms of prediction performance, we observe that model averaging offers the best overall performance on both the estimation and validation data.
+However, once we move outside that area and look at out-of-distribution prediction, the model averaging approach increases the weight for the models with behavioural underpinnings, sometimes substantially so.
+In terms of prediction performance, we observe that model averaging offers the best overall performance on both the estimation and test data.
 Crucially, this advantage is at its largest when making out-of-distribution predictions, in our case, predicting mode choice for trips whose distance is outside the range covered by the estimation data.
 There are, of course, some opportunities for further refinement.
-First, in our implementation, we used the same data to estimate the sub-models and the model averaging.
-In case of overfitting by one of the sub-models (an issue that did not arise in our case), this could create the risk that model averaging assigns disproportional weight to that model, which would jeopardise the benefits of the approach.
-In studies with larger sample sizes, a further splitting of the data could be considered.
-Second, in applying our proposed model averaging approach, an analyst needs to define what distribution means.
-In our work, we have focussed on the obvious example of trip distance.
+The key one relates to the fact that, in applying our proposed model averaging approach, an analyst needs to define what distribution means.
+In our work, we have focused on the obvious example of trip distance, for which there were pragmatic reasons, but also given that trip distance embodies key information in relation to similarity between trips.
 A core opportunity for future work is thus the investigation of out-of-distribution performance considering attributes other than trip distance alone, for example, going outside the range of socio-economic characteristics covered by the estimation data, or combinations of socio-economic characteristics not covered in the estimation data.
 Out-of-distribution could also take on a multivariate nature, thus covering multiple trip characteristics at the same time.
-Finally, we used MLP for the model averaging structure.
-Other meta-models, such as hybrid architectures approaches like TasteNetMNL (cf.
-Han et al., 2022), are an interesting avenue for future research.
+This, however, then also necessitates care by the analyst in defining what constitutes within-distribution and out-of-distribution.
+Finally, we employed an MLP as the meta-model within the model averaging framework.
+The literature on ensemble and stacking architectures in machine learning offers a broad range of methodological extensions that could further enhance such meta-modelling approaches (e.g.
+Shaikh et al., 2024).
+Future research could, for instance, investigate deep stacking networks, Bayesian ensemble formulations, or ensemble distillation methods to improve model robustness and generalisation, particularly in data-scarce settings.
+Furthermore, in cases where a larger number of explanatory variables are available and one wishes to understand their importance, more structured meta-models, such as a logistic regression model or hybrid models, such as TasteNet Han et al., 2022 can be explored.
 Acknowledgements

### v3

@@
 out-of-distribution prediction, and whether better predictions can be obtained by combining insights from the different models.
+We focus on trip distance as a key example of a variable where the application context might go beyond the estimation data.
 Across two case studies, we show that while data-driven approaches excel in predicting mode choice for trips within the distance bands used in estimation, beyond that range, the picture is fuzzy.
-To leverage the relative advantages of the different model families and capitalise on the notion that multiple ‘weak’ models can result in more robust models, we put forward the use of a model averaging approach that allocates weights to different model families as a function of the distance between the characteristics of the trip for which predictions are made, and those used in model estimation - in our case, we use trip distance as the sole characteristic for this.
+To leverage the relative advantages of the different model families and capitalise on the notion that multiple ‘weak’ models can result in more robust models, we put forward the use of a model averaging approach that allocates weights to different model families as a function of the distance between the characteristics of the trip for which predictions are made, and those used in model estimation.
 Overall, we see that the model averaging approach gives larger weight to models with stronger behavioural or econometric underpinnings the more we move outside the interval of trip distances covered in estimation.
@@
 Inference models and forecasting models are evaluated differently by analysts.
-When building an inference model, an analyst is keen to ensure that the model generates behaviourally plausible insights into explanatory factors and their relative impacts.
-When building a forecasting model, the focus is much more on a model’s ability to generate accurate predictions of behaviour.
-Importantly, in forecasting, we predict the behaviour in new settings, meaning that the models have to generalise out-of-sample.
-As some of the settings we wish to make predictions for may be outside the range of level-of-service values that the models have been estimated on, it is important for forecasting models to also generalise out-of-distribution.
+When building an inference model, the focus is on ensuring that the model generates behaviourally plausible insights into explanatory factors and their relative impacts.
+In contrast, when building a forecasting model, the emphasis is on the model’s ability to generate accurate predictions of behaviour.
+Importantly, forecasting involves predicting behaviour in new settings, meaning that the models have to generalise out-of-sample.
+This may relate, for example, to specific combinations of time and cost that were not observed in the estimation data.
+In addition, some prediction settings may lie outside the range of attribute values on which the models were estimated.
+This could involve higher costs, longer travel times, or, more intuitively, trips over longer distances [^1] We do not consider issues of spatial or temporal transferability (see e.g.
+Fox et al., 2014), nor the introduction of new modes or attributes.
+Instead, we focus on situations where attribute levels, or combinations thereof, fall outside the distribution observed in the estimation data.
+Our analysis is further restricted to continuous attributes, for which the impact of new attribute levels can be evaluated based on the estimated model parameters..
+In such cases, it is important for forecasting models to also generalise out-of-distribution.
 Travel behaviour analysts are increasingly interested in a more diverse set of models, covering not just econometric structures, but also models from mathematical psychology (e.g.
-Hancock et al., 2021, 2020a), and increasingly machine learning techniques (e.g.
+Hancock et al., 2021, 2020b), and increasingly machine learning techniques (e.g.
 Alwosheel et al., 2021; Wong and Farooq, 2021; Sfeir et al., 2022; Yan et al., 2025; Wang et al., 2020).
-This is especially true given the growing reliance on highly granualar data on travel behaviour, for example that collected using GPS tracking (e.g.
+This is especially true given the growing reliance on highly granular data on travel behaviour, for example that collected using GPS tracking (e.g.
 Liu et al., 2024; Tsoleridis et al., 2025).
@@
 Of course, the entire purpose of building a forecasting model is not to replicate choices that have already been made, but to make predictions for new settings.
-Despite the growing interest in machine learning approaches, the dominant - although not necessarily evidenced-based - view is that forecasting models with a solid behavioural underpinning, such as Random Utility Maximisation (RUM) based discrete choice models, are better equipped to predict behaviour under new settings than are models with a weak or no behaviour underpinning, such as e.g.
+Despite the growing interest in machine learning approaches, the dominant, although not necessarily evidenced-based, view is that forecasting models with a solid behavioural underpinning, such as Random Utility Maximisation (RUM) based discrete choice models, are better equipped to predict behaviour under new settings than are models with a weak or no behaviour underpinning, such as e.g.
 machine learning models (cf.
@@
 the model fit on the empirical data, in the process of developing forecasting models.
-What is clear is that good model performance is insufficient to establish that a model will make good out-of-distribution predictions.
-For that, the model must have captured underlying relationships that generalise beyond the observed data.
-The choice modelling literature has at times recognised that more advanced models that offer a better fit on the empirical data do not necessarily lead to better forecasts (see e.g.
-Fox et al., 2014).
-However, the question of how to develop forecasting models considering the model’s performance and behavioural underpinning is still open.
-On the one hand, the fact that model parameters are estimated by maximising the model performance on the empirical data, i.e.
-how well they replicate current choices, is somewhat at odds with the aim of forecasting, i.e.
-to generalise out-of-distribution.
-On the other hand, it is intuitive that the performance of a model on the empirical data still pertains to relevant information on its ability to generalise out-of-distribution.
-The aim of this paper is twofold.
+Performance on empirical (estimation) data is often taken as an indication of a model’s ability to generalise, despite evidence that more flexible models that achieve better in-sample fit do not necessarily lead to improved forecasts (see e.g.
+Fox et al., 2014; Hess and van Cranenburgh, 2025).
+Moreover, when behavioural patterns in out-of-distribution settings differ from those observed in the estimation data, a good in-sample fit alone may not translate into strong out-of-distribution performance.
+In such cases, models with a stronger behavioural underpinning may be better suited to capturing underlying decision mechanisms and could therefore generalise more robustly to new scenarios.
+The contribution of this paper is twofold.
 First, we aim to shed light on the prediction performance of forecasting models with varying levels of behavioural underpinning as a function of the “distance” between the training data and forecasting scenarios for models.
-Second, we aim to develop an approach that improves prediction performance in forecasts through a model averaging approach that assigns different weights to different models depending on this “distance".
-In our work, we use trip distance as the key variable, but the approach works with any variable that goes out of distribution, not only geographic distance.
+Second, we develop an approach to improve prediction performance through model averaging, where different weights are assigned to different models depending on this “distance".
+As highlighted above, there are many ways in which prediction settings could go outside the (multi-dimensional attribute) distribution covered by the estimation data.
+The focus of our work is on values of continuous attributes.
+We specifically use trip distance for this.
+It is a natural way of defining the context of a trip, and other key variables such as travel times and cost would be correlated with distance, thus also going out-of-distribution for trips that are longer or shorter than those included in the estimation data.
+There are also policy-relevant motivations for using distance as an example.
+In particular, forecasting behaviour beyond the range of observed trip distances is of interest in settings where long-distance trips are underrepresented in estimation data.
+While we use trip distance as the primary variable, the proposed approach is conceptually general and could be applied to other attributes that extend beyond the range observed in the estimation data.
+Assessing its performance in such settings is left for future work.
 We employ a neural network for the model averaging.
@@
 In model averaging (see e.g.
-Hancock et al., 2020b), which can be viewed as a form of stacking, first introduced in the seminal work of Wolpert (Wolpert, 1992), which established the conceptual basis for ensemble methods combining multiple predictive models.
+Hancock et al., 2020a), which can be viewed as a form of stacking, first introduced in the seminal work of Wolpert (Wolpert, 1992), which established the conceptual basis for ensemble methods combining multiple predictive models.
 Specifically, we first estimate $M$ different models on the data, with model $m$ giving a likelihood $L_{n,t,m}\left(\Omega_{m}\right)$ for the choice observed by decision maker $n$ in choice situation $t$, using a vector of estimated model parameters $\Omega_{m}$.
@@
 2.2 Model averaging for out-of-distribution prediction
-Model averaging typically estimates each model on the entire data and then computes the weights by considering how well each model fits for each of the data points, estimating such weights at the person or observation level.
-This could, for example, include decision-maker characteristics such as age, gender and income and trip characteristics such as purpose and trip distance.
-In the present paper, we focus on trip distance in particular.
-This is to some extent a pragmatic choice given the proof-of-concept nature of our work.
-It allows us to very easily split the data and define what constitutes within-distribution and out-of-distribution, with a view to demonstrating the performance of the model averaging approach.
-In addition however, trip distance has a number of appealing characteristics in this context.
-First, distance is not typically used as an explanatory variable in the mode choice models themselves, which instead rely on the closely correlated variables such as time and cost.
-Second, distance can explain much of the similarity between trips, along multiple dimensions at the same time.
-This way, the model averaging structure can learn which model works well for what trip distance, without having to rely on multiple covariates, given that this single variable condenses other information into a single factor.
+Model averaging typically involves estimating each model on the full dataset and then computing weights based on how well each model fits for each of the data points, estimating such weights at the person or observation level.
+These weights can depend on decision-maker characteristics such as age, gender, and income, as well as trip characteristics such as purpose and trip distance.
+In the present paper, as outlined in the introduction, we focus specifically on trip distance.
+This is partly a pragmatic choice, reflecting the proof-of-concept nature of our work.
+It allows for a straightforward definition of in-distribution and out-of-distribution settings through a simple data split, facilitating a clear assessment of the proposed model averaging approach.
+It should be noted that this definition of out-of-distribution prediction may favour models with relatively high bias and low variance, as such models often extrapolate more smoothly beyond the observed range (Bishop and Nasrabadi, 2006).
+Trip distance also has several appealing characteristics in this context.
+First, it is typically not included directly as an explanatory variable in the mode choice models, which instead rely on closely correlated variables such as time and cost.
+Second, distance provides a parsimonious way of capturing variation across trips, as it is correlated with multiple underlying attributes.
+This allows the model averaging structure to learn how model performance varies with trip context using a single, interpretable variable, rather than relying on a higher-dimensional set of covariates.
 Training a model averaging structure where journey distance, say $d_{n,t}$ for person $n$ and trip $t$, is used as a key characteristic in Equation 4 would allow the model to learn how different sub-models are better suited for predicting trips of specific distances within the bounds of the estimation data.
@@
 In the estimation of the individual sub-models, we then use only a subset of the range of the available data, specifically those observations where $d_{n,t}\in D_{train}$.
-The model averaging structure on the other hand will be estimated on the entire data, i.e.
+The model averaging structure, on the other hand, will be estimated on the entire data, i.e.
 still maximising e.g.
@@
 We specifically use the GPS tracking component of this survey, where the data used for this work corresponds to the observed mode choice behaviour.
+Respondents were tracked over a period of two weeks, with most trips taking place in Yorkshire.
 After extensive data cleaning and data enrichment (Tsoleridis et al., 2022), 10,990 trips made by 415 individuals remained.
 For each trip, individuals travelled by one of six modes: car (48%), bus (14%), rail (5%), taxi (3%), cycling (4%) or walking (26%).
-The data covers a very wide range of trip distances, going from just under 100 metres to just under $106$ kilometres, with a mean distance of $8.2$ kilometres and a median distance of $4.5$ kilometres.
+The data covers a wide range of trip distances, going from just under 100 metres to just under $106$ kilometres, with a mean distance of $8.2$ kilometres and a median distance of $4.5$ kilometres.
+It thus covers both urban trips as well as inter-city trips, for example between Leeds and York.
 The attributes of the alternatives used in the models for the present paper include in-vehicle travel time, out-of-vehicle travel time, and travel cost.
@@
 This dataset contains four alternatives: walking (18%), cycling (3%), public transport (35%, grouping together bus and rail) and driving (44%, grouping together private car and taxi rides).
-The data again covers a very wide range of trip distances, going from just under 100 metres to just over $40$ kilometres, with a mean distance of $4.6$ kilometres and a median distance of $2.8$ kilometres.
-Attributes of the alternatives used in the models include in-vehicle travel time, out-of-vehicle travel time, interchanges, and travel cost, along with two socio-demographic variables, namely car ownership and driving license status.
+The data again covers a wide range of trip distances, going from just under 100 metres to just over $40$ kilometres, with a mean distance of $4.6$ kilometres and a median distance of $2.8$ kilometres.
+This dataset covers all of London, including short-distance trips as well as journeys from one side of London to the other.
+Attributes of the alternatives used in the models include in-vehicle travel time, out-of-vehicle travel time, interchanges, and travel cost, along with two sociodemographic variables, namely car ownership and driving licence status.
 3.2 Estimation and test data
@@
 For the purposes of our analysis, we set $d_{a,MA}$ and $d_{b,MA}$ to be at the 10th and 90th percentiles of the sample distance distribution $D_{data}$, while $d_{a,sub}$ and $d_{b,sub}$ are set at the 20th and 80th percentiles of the sample distance distribution $D_{data}$.
+As outlined above, the shortest 10% and longest 10% of trips were retained for out-of-distribution testing.
+In both datasets, this implies that the out-of-distribution portion corresponds to the tails of the distance distribution.
+The upper tail is particularly extensive, covering trips between 19.52 km and 105.9 km in the DECISIONS data, and between 11.22 km and 40.1 km in the LPMC data.
+This represents over 80% of the distance range in the DECISIONS data and over 70% in the LPMC data.
+This creates a meaningful out-of-distribution setting, with inter-city trips in the DECISIONS data and cross-London trips in the LPMC data appearing exclusively in the test set, meaning that the contexts are structurally different from the estimation data.
 In addition to these divisions, we further retain $20\%$ of any trips, independent of their distance, for out-of-sample test.
-The splitting of the sample into estimation, test (and additionally validation data for MLP and XGB) was done at the person/household level, so as to avoid data leakage.
+The splitting of the sample into estimation, test (and additionally validation data for MLP and XGB) was done at the person/household level, to avoid data leakage.
 In other words, each person only contributed data to either the estimate data or the test data.
@@
 The resulting division of the data is explained in Figure 1, with the resulting sample sizes shown Table 1.
-In practical work, analysts would be unlikely to not include all deciles (intervals of 10% width) in the MA estimation work - we made these decisions with a view to validating the model idea.
+In practical work, analysts would be unlikely to not include all deciles (intervals of 10% width) in the MA estimation work; we made these decisions with a view to validating the model idea.
 Similarly, a real-world application would not set aside the test data, and thus also avoid deleting observations in the estimation data for trips that are in segments 1 and 10 (see the earlier point about data leakage).
@@
 Individual MA Individual MA
-out of distribution for MA (below, i.e. $d_{min}-d_{a,MA}$ <1.33 0 0 223 <0.75 0 0 1,581
+out of distribution for MA (below, i.e. $d_{min}-d_{a,MA}$) <1.33 0 0 223 <0.75 0 0 1,581
 out-of-distribution for sub-models (below, i.e. $d_{a,MA}$ to $d_{a,sub}$) 1.33 to 1.95 0 836 261 0.75 to 1.11 0 6,446 1,642
@@
 The following individual models are used in our analysis, combining structures from choice modelling, mathematical psychology and machine learning.
-Many other model types could have been considered [^1] A notable omission in our list is mixture models such as mixed logit or latent class - any advantages that such models have in estimation largely/completely disappear in prediction as the random heterogeneity needs to be averaged out.
-The only exception comes with the use of posterior distributions, which are not applicable in the context of out-of-sample prediction.
-- we limited ourselves to this set with a view to demonstrating the potential benefit of our proposed approach.
+Many other model types could have been considered [^2] A notable omission in our list is mixture models such as mixed logit or latent class.
+Any advantages that such models have in estimation largely/completely disappear in prediction as the random heterogeneity needs to be averaged out.
+The only exception comes with the use of posterior distributions, which are not applicable in the context of out-of-sample prediction., where we limited ourselves to this set with a view to demonstrating the potential benefit of our proposed approach.
 Nested Logit (NL):
@@
 The MLP is implemented in PyTorch (Pedregosa et al., 2011) using a custom to handle the availability conditions, while the XGBoost model is implemented directly with the xgboost package (Chen and Guestrin, 2016).
-For both models, a hyperparameter grid search was conducted to optimise performance.
+For the two data-driven models, a hyperparameter grid search was conducted to optimise performance.
 We used a 5-fold Grouped cross-validation procedure to account for the panel structure of the data (see Table 2).
-The hyperparameter ranges for the XGB model were informed by the results of Hillel (2021).
-The number of hidden layers and nodes per layer influences the model’s capacity to learn complex patterns, while the batch size affects the stability and efficiency of the gradient updates.
+For MLP, the number of hidden layers and nodes per layer influences the model’s capacity to learn complex patterns, while the batch size affects the stability and efficiency of the gradient updates.
 The learning rate controls how quickly the model updates its weights during training, while the L2 regularisation helps prevent overfitting by penalising large weight values.
@@
 The best-performing configuration found was two hidden layers with 30 nodes each, a learning rate of 0.001, a batch size of 250 and an L2 regularisation strength of 0.
-Because each training run of an MLP can yield slightly different outcomes, we trained the model 100 times using the best-performing configuration and averaged its predictions to account for this variability.
-For the XGBoost model, the hyperparameter search included the maximum tree depth, the number of estimators, the minimum loss reduction, the minimum child weight, the subsample fraction, the fraction of features considered per split, regularisation strengths (L1 and L2), and the maximum delta step.
-The maximum tree depth controls the complexity of each tree, with deeper trees allowing more detailed partitions at the risk of overfitting.
-The number of estimators determines how many boosting rounds are performed, thereby influencing the overall capacity of the model.
-The minimum loss reduction acts as a pruning mechanism by preventing splits that do not provide sufficient improvement to the objective function, while the minimum child weight enforces a lower bound on the sum of instance weights required in a leaf, discouraging overly specific partitions.
+An important point to highlight is variability in results.
+Each training run of an MLP can yield slightly different outcomes.
+An analyst then needs to make a decision on how many training runs to use, and which of the results to retain.
+We trained the model 100 times using the configuration described above.
+If our focus at this stage was purely on prediction performance, we would concentrate on the best fitting models out of these runs.
+However, our objective here is also to consider diversity across model instances, and capture variability arising from different local optima and promote ensemble diversity.
+We thus and averaged the predictions across these 100 runs to account for this variability.
+The hyperparameter ranges for the XGB model were informed by the results of Hillel (2021).
+The hyperparameter search included the maximum tree depth, the subsample fraction, the fraction of features considered per split, regularisation strengths (L1 and L2), and the maximum delta step.
+Furthermore, we set the learning rate to 0.1 and used the default values for all other hyperparameters.
+During training, early stopping was used to control the number of boosting rounds (with a maximum set to 3000).
+Thus, training was stopped if adding more trees did not further improve performance (on the validation set) after 10 boosting rounds.
+About the hyperparameter: the maximum tree depth controls the complexity of each tree, with deeper trees allowing more detailed partitions at the risk of overfitting.
 The subsample fraction regulates the proportion of training samples used for each boosting round, and the feature subsampling parameter (colsample by tree) increases model robustness by introducing randomness at the feature level.
@@
 Finally, the maximum delta step restricts the weight update size, improving convergence stability in cases of unbalanced data.
-Unlike the MLPs, the optimal hyperparameters for the XGBoost model were not identical across the datasets.
-Both the London and Decision data shared a fixed learning rate of 0.01, a minimum loss reduction of 0.6, a minimum child weight of 20, and a subsample fraction of 0.7.
-However, the London configuration used a feature subsampling rate of 0.75, a maximum tree depth of 6, 1,200 boosting rounds, and regularisation strengths of $L_{1}=20$ and $L_{2}=40$, with a maximum delta step of 6.
-In contrast, the Decision configuration relied on full feature sampling (1.0), a deeper maximum tree depth of 9, 800 boosting rounds, and lighter $L_{1}$ regularisation of 10 (with $L_{2}=40$), combined with a smaller maximum delta step of 4.
+The optimal hyperparameters for LPMC were: Max_tree_depth = 4, Subsample_fraction = 0.8, Colsample_by_tree = 0.8, $L_{1}$ reg.
+= 10, $L_{2}$ reg.
+= 1, and max_delta_set = 1.
+The optimal hyperparameters for DECISIONS were: Max_tree_depth = 2, Subsample_fraction = 0.4, Colsample_by_tree = 0.6, $L_{1}$ reg.
+= 1, $L_{2}$ reg.
+= 20, and max_delta_set = 2.
 As with the MLP model, we repeated the training 100 times using the optimal configuration and averaged their predictions to account for variability in the model training.
@@
 Hidden layers: $\{(10,5),(10,10),$ (20,20),(30,30)}\;(20,20),(30,30)\} Batch size: $\{250,500,1000\}$ Learning rate: $\{0.01,0.001,0.0001\}$ $L_{2}$ regularisation: $\{0,0.1,0.5,1\}$ Max.
-tree depth: $\{6,9\}$ Number of estimators: $\{800,1200\}$ Min.
-loss reduction: $\{0.2,0.6\}$ Min.
-child weight: $\{10,20\}$ Subsample fraction: $\{0.7,0.8\}$ Colsample by tree: $\{0.75,1.0\}$ $L_{1}$ regularisation: $\{10,20,40\}$ $L_{2}$ regularisation: $\{10,20,40\}$ Max.
-delta step: $\{4,6\}$
+tree depth: $\{2,3,4,5\}$ Subsample fraction: $\{0.4,0.6,0.8\}$ Colsample by tree: $\{0.4,0.6,0.8\}$ $L_{1}$ regularisation: $\{1,5,10,20\}$ $L_{2}$ regularisation: $\{1,5,10,20\}$ Max.
+delta step: $\{0,1,2,4\}$
 3.4 Model averaging: specification and estimation
-Similar to the MLP models implemented for the training on the data sets (see Section 3.3), we implemented the MLP for the model averaging in PyTorch Paszke et al.
-(2019) too.
-This is needed because the loss function (see equation (2)) requires a custom implementation, which is not possible in e.g.
-SciKit-learn.
+Similar to the MLP models used for training on the datasets (see Section 3.3), the MLP used for model averaging was implemented in PyTorch Paszke et al.
+(2019).
+This implementation is required because the loss function (see Equation (2)) involves a custom specification that cannot be readily implemented in, for example, Scikit-learn.
 As explanatory features, we included three distance-based variables: the raw trip distance, its logarithmic transformation, and its squared term.
-While in theory the linear distance alone should suffice, we found that adding the log- and squared transformations provided a small but consistent improvement in model performance.
-To tune the hyperparameters, we performed a grid search using a simple 75%–25% train–test split.
-Ideally, a dedicated validation set would have been used for this purpose—separate from the data employed in training the individual models, see Hastie (2009), who argues that not doing so might result in the meta-model assigning unfairly high weights to models with higher complexity.
-However, we deliberately refrained from further partitioning the data to avoid diminishing the already limited estimation sample and to ensure that the test set remained completely untouched for the final performance evaluation.
-The following parameter spaces are used: batch size (64, 128, 256), number of hidden layers and nodes (two or three hidden layers with each 10, 20, 40, or 60 nodes), learning rate (0.001, 0.005, 0.01) and L2 regularisation (0, 0.0001, 0.001).
-For the DECISIONS data set, the best-performing model has two hidden layers with 20 nodes each, a batch size of 128, a learning rate of 0.005, and an L2 regularisation strength of 0.001.
-For the LPMC data set, the best-performing model has two hidden layers with 60 nodes each, a batch size of 64, a learning rate of 0.001, and an L2 regularisation strength of 0.0001.
-We repeated the training 100 times using the optimal configuration.
-As we aim for high prediction performance, we retained only the top 20 per cent of models [^2] This selection was based on the log-likelihood.
+While in principle the linear distance alone should suffice, we found that including the logarithmic and squared transformations resulted in a small but consistent improvement in predictive performance.
+Figure 2 shows the structure for our model averaging approach.
+As shown in the figure, the model averaging MLP only takes the trip distance features as input to predict the weights given to each sub-model.
+The model-averaging probability is computed as a weighted sum, combining the predictions for the trip by each sub-model and the weights assigned to each sub-model by the model-averaging MLP.
+Hyperparameters were tuned using a grid search combined with five-fold cross-validation, where folds were defined at the individual (household_id) level to avoid leakage across observations from the same individual.
+The following parameter spaces were considered: batch size (50, 100, 250), one hidden layer with 10, 20, or 30 nodes, learning rate (0.01, 0.001, 0.0001), and L2 regularisation strength (0, 0.01, 0.05).
+For the DECISIONS dataset, the best-performing model has a hidden layer with 10 nodes, a batch size of 50, a learning rate of 0.01, and an L2 regularisation strength of 0.
+For the LPMC dataset, the best-performing model has a hidden layer with 10 nodes, a batch size of 250, a learning rate of 0.01, and an L2 regularisation strength of 0.
+Using the optimal hyperparameters, we repeated the training 100 times.
+In contrast with the training of the individual models, our objective is now high predictive performance.
+We therefore retained only the top 20 per cent of models [^3] This selection was based on the log-likelihood.
 For this, we used the full data, i.e.
 training and validation data, but excluded the test data., discarding those that became trapped in poor local maxima.
-This approach, thus, differs slightly from the one used for the training of the data-driven models, where we also repeated training 100 times but averaged predictions across all models to encourage diversity between models.
-Figure 2 shows the structure for our model averaging approach.
-As shown in the figure, the model averaging MLP only takes the trip distance as input to predict the weights given to each sub-model.
-The model-averaging probability is computed as a weighted sum, combining the predictions for the trip by each sub-model and the weights assigned to each sub-model by the model-averaging MLP.
+This approach differs slightly from the one used for training the data-driven models, where we also repeated training 100 times but averaged predictions across all models to encourage diversity, while diversity is now captured by the different models entering the ensemble.
+It should be noted that the data used to train the meta-model are the same observations used to train the base models (for distance segments 3–8).
+Hastie et al.
+(2009) argue that using the same data for training both the base models and the meta-model may lead the meta-model to assign disproportionately high weights to more complex models, as their in-sample predictions can appear overly optimistic.
+To mitigate this issue, they propose a stacking procedure in which the meta-model is trained on predictions generated from data not used in training the base models (often obtained via cross-validated predictions).
+To further facilitate comparison with the machine learning literature on stacked generalisation, Table 3 provides a mapping between the terminology used in this paper and the standard terminology used in that literature.
+While we recognise the concern about using the same for both the base models and meta-model, we deliberately refrained from implementing a stacking procedure in the main analysis in order to keep the modelling pipeline tractable.
+The overall framework already involves a large number of models, data partitions, hyperparameter tunings, and aggregation steps, and introducing a full cross-validated stacking procedure would substantially increase the computational and methodological complexity.
+Nevertheless, to assess whether this modelling choice affects our results, we conducted an additional analysis on the larger LPMC dataset in which the meta-model is trained on predictions generated from data not used to train the base models.
+The results of this analysis, reported in Appendix A, indicate that the impact on predictive performance is negligible, suggesting that the simpler procedure used in the main analysis does not materially affect the conclusions, at least in our application.
 Figure 2: Model averaging structure
+Table 3: Terminology mapping between this study and the stacking literature
+Current paper Stacking literature
+Sub-model / individual model Base learner (level-0 model)
+Model averaging model Meta-learner (level-1 model)
+Model averaging weights Combination weights / stacking weights
+Predicted probabilities (per model) Base learner predictions
+Ensemble / model averaging approach Stacked generalisation / stacking
+Distance-based features Meta-features
 3.5 Empirical results
@@
 MLP -1,614.22 -22,154.49
-XGB -1,650.34 -21,628.70
+XGB -1,691.79 -21,031.13
 As a next step, Figure 3 looks at the performance of these models across the different distance segments obtained by splitting the data into deciles of the distance distribution.
@@
 segments 3-8).
-Across the distance segments, there are some where XGB outperforms MLP in the DECISIONS data, or MLP outperforms XGB in LPMC (reversing the overall trend), while the overall trend between DFT and NL/MNL is also reversed in some cases.
+Across the distance segments, there are no cases where XGB outperforms MLP in the DECISIONS data, or MLP outperforms XGB in LPMC (which would reverse the overall trend), but differences change in size.
+Meanwhile, the overall trend between DFT and NL/MNL is reversed in some cases.
 These findings already show that different models perform differently well across trip distances.
-Interestingly, while the LPMC results show that mode choice is easier to predict for longer trips (across segments 3-8), no clear trend emerges for the DECISIONS data.
+Interestingly, while the LPMC results suggest that mode choice is easier to predict for longer trips (across segments 3-8), no clear trend emerges for the DECISIONS data.
 We next turn to segments 2 and 9 in the left two graphs, giving us initial insights into out-of-distribution performance, remembering that the values shown here are the predicted probabilities for the chosen alternative, using the model parameters obtained when estimating the models on segments 3 to 8.
@@
 We first focus on segments 3-8 with a view to understanding potential overfitting to the estimation data.
-For the DECISIONS data, we see some evidence of improved out-of-sample (vs within-sample) performance for all models, with an average gain in per observation LL of $0.096$ for NL, $0.095$ for DFT, $0.062$ for MLP, and $0.054$ for XGB.
-The same happens for LPMC models, the same is observed, with improvements by $0.12$ for NL and DFT, $0.11$ for MLP and $0.094$ for XGB.
+For the DECISIONS data, we see some evidence of improved out-of-sample (vs within-sample) performance for all models, with an average gain in per observation LL of $0.096$ for NL, $0.095$ for DFT, $0.062$ for MLP, and $0.065$ for XGB.
+The same happens for LPMC models, with improvements by $0.12$ for NL and DFT, $0.11$ for MLP and $0.081$ for XGB.
 Overall, these results do not suggest any overfitting, at least within the distribution of distances covered by the estimation data.
-The findings for segments 2 and 9 on the test data largely confirm those we saw in the two graphs on the left in Figure 3, except for MLP and XGB performing more poorly in the test data for segment 9.
+The findings for segments 2 and 9 on the test data largely confirm those we saw in the two graphs on the left in Figure 3, except for MLP and XGB performing more poorly in the test data for segment 9 for the DECISIONS data.
 We finally turn to segments 1 and 10.
@@
 For the longest trips (segment 10), the trend is less clear.
-In the DECISIONS data, MLP performs comparatively well.
-In the LPMC data, both data-driven models do comparatively well, while MNL suffers a big drop.
+In the DECISIONS data, MLP and XGB perform comparatively well.
+In the LPMC data, both data-driven models again do comparatively well, while MNL suffers a big drop.
 3.5.2 Model averaging results
@@
 Model DECISIONS LPMC
-NL/MNL -3,134.33 (11.15%) -40,210.60 (4.49%)
-DFT -3,061.83 (9.04%) -40,186.25 (4.43%)
-MLP -2,846.66 (2.17%) -38,966.05 (1.44%)
-XGB -2,999.89 (7.16%) -38,882.09 (1.23%)
-MA -2,784.99 -38,404.69
+NL/MNL -3,134.33 (10.71%) -40,210.60 (6.15%)
+DFT -3,061.83 (8.60%) -40,186.25 (6.10%)
+MLP -2,846.66 (1.69%) -38,966.05 (3.16%)
+XGB -2,998.23 (6.66%) -38,182.65 (1.17%)
+MA -2,798.65 -37,736.03
 The inclusion of segments 2 and 9 allows the model averaging structure to learn about out-of-distribution performance and hence guide the process of linking the weights given to individual models to the difference between the distance for that trip and the data used in estimation (i.e.
@@
 For both datasets, we see that the MA process gives almost the entire weight to MLP and XGB for any trips with a distance contained within $D_{train,sub}$, i.e.
-distances on which the individual sub-models were estimated.
-In both datasets, XGB gets higher weight for short trips within that part of the distance distribution compared to long trips, while for longer trips, almost the entire weight goes to MLP for DECISIONS, while it also increases for LPMC.
-These findings are in line with the pattern of prediction performance obtained for individual models, as shown on the left hand side in Figure 3.
+distances on which the individual sub-models were estimated (segments 3-8).
+In the DECISIONS data, the weight of XGB increases on shorter trips within that part of the distance distribution compared to long trips, while for longer trips, almost the entire weight goes to MLP for DECISIONS.
+For LPMC, all the weight goes to XGB.
+These findings are in line with the pattern of prediction performance obtained for individual models, as shown on the left-hand side in Figure 3.
 Once we go outside the $D_{train,sub}$ interval, the pattern changes.
@@
 segments 1&2), the weight for behavioural models increases.
-In the DECISIONS data, only a small share goes to NL, with quite equal weight for the other three models.
-On the other hand, in LPMC, MNL essentially obtains all the weight.
+In the DECISIONS data, only a small share goes to NL and DFT.
+On the other hand, in LPMC, MNL obtains over 80% of the weight.
 For longer trips, i.e.
-above $d_{b,sub}$, NL gets the vast majority of the weight in the DECISIONS data, with the rest going to DFT.
+above $d_{b,sub}$, NL gets the vast majority of the weight in the DECISIONS data, with DFT getting more than MLP or XGB.
 For LPMC, MLP retains a large weight, but this decreases while the weight for MNL and DFT increases with distance.
@@
 Table 4.
-For both datasets, the gains by model averaging over the best performing sub-model are larger on the test data than on the estimation data.
+For both datasets, the gains by model averaging over the best-performing sub-model are larger on the test data than on the estimation data.
 Table 6: Log-Likelihood for model averaging: individual components and overall structure on test data (covering distance segments 1-10), with gains by MA over individual models shown in brackets
 Model DECISIONS LPMC
-NL/MNL -825.88 (6.12%) -11,764.56 (5.30%)
-DFT -805.78 (3.78%) -11,426.72 (2.50%)
-MLP -804.91 (3.68%) -11,448.70 (2.68%)
-XGB -829.64 (6.55%) -12,082.56 (7.79%)
-MA -775.32 -11,141.55
+NL/MNL -825.88 (7.14%) -11,764.56 (5.52%)
+DFT -805.78 (4.82%) -11,426.72 (2.72%)
+MLP -804.91 (4.72%) -11,448.70 (2.91%)
+XGB -806.62 (4.92%) -12,037.51 (7.66%)
+MA -766.88 -11,115.43
 There are two possible influences that could drive the differences between Table 5 and Table 6.
@@
 To support this discussion, Figure 5 shows the per observation log-likelihoods in the different segments for estimation and test.
-We first see that, in estimation, model averaging outperforms (or equals) the four sub-models in 4 out of 8 estimation segments for the DECISIONS data, and 6 out of 8 for the LPMC data.
-When we turn to the test data, model averaging performs best (or joint best) in 2 out of 10 segments for the DECISIONS data, and 5 out of 10 segments for the LPMC data.
-These per segment statistics go hand in hand with the best overall LL obtained by model averaging as reported in Table 4 and Table 5.
+We first see that, in estimation, model averaging outperforms (or equals) the four sub-models in 6 out of 8 estimation segments for the DECISIONS data, and 7 out of 8 for the LPMC data.
+When we turn to the test data, model averaging performs best (or joint best) in 4 out of 10 segments for the DECISIONS data, and 5 out of 10 segments for the LPMC data.
+These per-segment statistics go hand in hand with the best overall LL obtained by model averaging as reported in Table 5 and Table 6.
 Figure 5: MA Estimation and test performance
-We earlier raised two possible reasons for why the advantages of MA might be larger on the test data than on the validation data.
+We earlier raised two possible reasons why the advantages of MA might be larger on the test data than on the validation data.
 In relation to over-fitting to the estimation data, we see that, for all four sub-models as well as for the model averaging structure, there are segments (out of 2-9) where the per observation LL increases and segments where it decreases.
 No clear pattern emerges in the comparison across models and segments.
-Turning to the second potential reason, namely out-of-distribution performance, we observe that for the DECISIONS data, model averaging is outperformed by NL in segment 1, but obtains by far the best performance in segment 10, where its combined performance on segments 1 and 10 is much higher than for any of the other models (-0.2582 vs -0.2944 for XGB).
+Turning to the second potential reason, namely out-of-distribution performance, we observe that for the DECISIONS data, model averaging is outperformed by NL in segment 1, but obtains by far the best performance in segment 10, where its combined performance on segments 1 and 10 is higher than for any of the other models (-0.248 vs -0.252 for XGB).
 For the LPMC data, a similar story emerges.
 Model averaging is outperformed slightly by NL in segment 1, but obtains by far the best performance in segment 10.
-Again, overall, MA performs the best of all models when looking at segments 1 and 10 together (-0.5793 vs -0.6025 for DFT).
+Again, overall, MA performs the best of all models when looking at segments 1 and 10 together (-0.577 vs -0.6025 for DFT).
 Overall, these findings lead to the conclusion that there is little difference across the models in terms of out-of-sample performance when looking at segments 2-9.
@@
 Crucially, this advantage is at its largest when making out-of-distribution predictions, in our case, predicting mode choice for trips whose distance is outside the range covered by the estimation data.
+While the analysis focuses on predictive performance, the observed differences across model classes can be interpreted in light of their underlying structure.
+In particular, models with a stronger behavioural underpinning may exhibit more stable performance in out-of-distribution settings because their parameters are constrained by behavioural theory and regularity conditions.
+This can limit the extent to which such models capture spurious patterns or context-specific correlations present in the estimation data.
+In contrast, more flexible, data-driven models may achieve superior in-sample fit by capturing patterns that do not generalise, but this flexibility can lead to reduced robustness when extrapolating beyond the observed data.
+We emphasise that this interpretation remains tentative and is not directly tested in this study, but it provides a plausible explanation for the observed performance differences and aligns with broader discussions on the trade-off between flexibility and generalisation.
+Our work relies on multiple separate models, and this clearly leads to increased complexity as well as doing nothing to address the black box nature of machine learning approaches.
+A contrasting approach is to make the trade-off between fit, generalisation, and interpretability more explicit at the estimation stage.
+Recent work has started to do so by proposing multi-objective estimation frameworks that jointly consider in-sample fit and out-of-sample predictive accuracy, thereby making it possible to explore Pareto-efficient specifications rather than selecting models on fit alone (Beeramoole et al., 2024).
+Related work has also shown that machine learning methods can be made more behaviourally credible by embedding behavioural regularity or interpretability constraints, which may help improve robustness when moving beyond the estimation domain (Feng et al., 2024).
+At the same time, the recent benchmarking literature makes clear that choice modellers are right to take the predictive strengths of machine learning seriously: across a very large empirical comparison, many ML models outperform traditional discrete choice models in prediction accuracy, even while important questions remain around transferability and uncertainty (Wang et al., 2024).
+It would therefore be unwise to disregard the strong in-sample performance that data-driven methods can deliver.
+Instead, our model averaging approach is motivated precisely by the idea of combining the strengths of both traditions: the behavioural interpretability and theoretical discipline of choice models, and the empirical flexibility and predictive power of machine learning methods.
 There are, of course, some opportunities for further refinement.
@@
 In our work, we have focused on the obvious example of trip distance, for which there were pragmatic reasons, but also given that trip distance embodies key information in relation to similarity between trips.
-A core opportunity for future work is thus the investigation of out-of-distribution performance considering attributes other than trip distance alone, for example, going outside the range of socio-economic characteristics covered by the estimation data, or combinations of socio-economic characteristics not covered in the estimation data.
+A core opportunity for future work is thus the investigation of out-of-distribution performance considering attributes besides trip distance alone, for example, going outside the range of socio-economic characteristics covered by the estimation data, or combinations of socio-economic characteristics not covered in the estimation data.
 Out-of-distribution could also take on a multivariate nature, thus covering multiple trip characteristics at the same time.
@@
 The authors are thankful to Thomas Hancock, Tim Hillel and Georges Sfeir for comments on an earlier draft.
+The comments of three anonymous referees have helped us to substantially improve the paper.
+Appendix A Cross-validated stacking LPMC
+To assess whether the absence of a separate validation set for model averaging affects the results, we conducted an additional analysis using the larger LPMC dataset in which a dedicated split is introduced.
+First, the base models (level-0 models) are trained exclusively on the estimation data.
+These trained models are then applied to a separate validation set that was not used during the training of the base models.
+The predictions generated by the base models on this validation set form the inputs for the model averaging stage.
+Next, the model averaging MLP (level-1 model) is trained using this validation data.
+To facilitate training and prevent overfitting, the validation set is further split into a val_train and val_test subset.
+The val_train subset is used to estimate the parameters of the model averaging MLP, while the val_test subset is used to monitor predictive performance during training.
+Training is stopped early when performance on the val_test subset deteriorates, implementing an early stopping criterion.
+This procedure ensures that the level-1 model is estimated using predictions obtained from data that were not used to train the base models, thereby providing a stricter separation between training stages.
+While this procedure is not identical to cross-validated stacking as typically implemented in the stacking literature, it closely mirrors the workflow used in the main analysis.
+As such, it provides a meaningful assessment of whether the absence of a dedicated validation set for model averaging affects the results reported in the main body of the paper.
+The resulting splitting is illustrated in Figure 6.
+Figure 6: Data splitting with validation
+Below, we present the main results to facilitate comparison with those reported in the paper.
+The fit of the model averaging approach on the test data (covering distance segments 1-10) remains largely unchanged, with a log-likelihood of −11,204.87, vs.
+-11,115.43 in the paper, implying a drop in LL by $0.8\%$.
+More importantly, Figure 7 shows broadly the same mixing pattern as a function of distance.
+One difference is that the allocation of weights across models is somewhat sharper and less gradual.
+In particular, within distance segments 3–8 virtually all weight is assigned to XGB, whereas outside this range XGB receives almost no probability mass.
+In segment 9, MLP attains the largest share of the probability, while segments 1 and 10 exhibit a mixture of MNL, DFT, and MLP.
+Figure 7: Heatmap of model performance across distance segments.
+A comparison of Figure 8 with the results reported in the paper reveals a largely similar pattern.
+The numerical values change only marginally.
+The only noteworthy difference is a modest increase in the mean log-likelihood achieved by model averaging in distance segment 1, which rises to 0.726-0.726 (vs.
+0.658-0.658 in the paper).
+The slightly sharper allocation may be related to the revised estimation procedure, where the level-0 models and the meta-model are trained on separate data partitions following Hastie et al.
+(2009).
+Training the meta-model on predictions that reflect out-of-sample performance can affect the allocation of weights across models.
+However, given the very modest changes observed, the results suggest that not using a separate validation set for model averaging has little impact on our overall findings.
+Figure 8: Heatmap of model performance across distance segments.